### v2

# Strong Linearizability without Compare&Swap: The Case of Bags

Abstract
Because strongly-linearizable objects provide stronger guarantees than linearizability, they serve as valuable building blocks for the design of concurrent data structures. Yet, many objects that have linearizable implementations from base objects weaker than compare&swap objects do not have strongly-linearizable implementations from the same base objects. We focus on one such object: the bag, a multiset from which processes can take unspecified elements.
We present the first lock-free, strongly-linearizable implementation of a bag from interfering objects (specifically, registers, and test&set objects). This may be surprising, since there are provably no such implementations of stacks or queues.
Since a bag can contain arbitrarily many elements, an unbounded amount of space must be used to implement it. Hence, it makes sense to also consider a bag with a bound on its capacity. However, like stacks and queues, a bag with capacity $b$ shared by more than $2b$ processes has no lock-free, strongly-linearizable implementation from interfering objects. If we further restrict a bounded bag so that only one process can insert into it, we are able to obtain a lock-free, strongly-linearizable implementation from $O(b+n)$ interfering objects, where $n$ is the number of processes.
Our goal is to understand the circumstances under which strongly-linearizable implementations of bags exist and, more generally, to understand the power of interfering objects.
Keywords: Strong-Linearizability, Bag, Concurrent Data Structures, Wait-Freedom, Lock-Freedom
1 Introduction
Concurrent data structures are often built using linearizable implementations of objects as if they were atomic objects. Although linearizability is the usual correctness condition for concurrent algorithms, there are scenarios where composing linearizable implementations does not preserve desired properties. Namely, an algorithm that uses atomic objects may lose some of its properties when these atomic objects are replaced by linearizable implementations. Specifically, linearizability does not preserve hyperproperties (properties of sets of executions), for example, probability distributions of events for randomized programs and security properties such as noninterference [10, 6]. This can be rectified by using strongly-linearizable implementations [10]. They guarantee that the linearization of a prefix of a concurrent execution is a prefix of the linearization of the whole execution. This means that the linearization of operations in an execution prefix cannot depend on later events in the execution.
Attiya, Castaneda, and Enea [3] observed that strongly-linearizable implementations of objects typically use compare&swap objects, which can be used to solve consensus among any number of processes. They provided strongly-linearizable implementations of some objects, such as a wait-free single-writer snapshot object and a wait-free max register from a fetch&add object, and a lock-free, readable, resettable test&set object and a lock-free, readable, fetch&increment object from an infinite array of test&set objects and registers. Both fetch&add and test&set objects are less powerful than a compare&swap object.
In this work, we aim to further explore the power of these well-studied building blocks in the context of strong linearizability. Specifically, we investigate strongly-linearizable implementations of a bag from these primitives. Establishing the existence of such implementations deepens our understanding of the circumstances under which these primitives can be used to achieve strong linearizability.
A bag is a multiset to which processes can insert elements and from which they can take unspecified elements. A concurrent bag is an abstraction of the interaction between producers and consumers and is commonly used to distribute tasks among processes for load balancing and improved scalability.
Queues and stacks have strongly-linearizable implementations from compare&swap objects [15, 21, 19]. A bag has a straightforward implementation from a stack or a queue. Therefore, a bag has a strongly-linearizable implementation from compare&swap objects. Although there is a wait-free, linearizable implementation of a stack from registers, test&set objects, and a readable fetch&increment object [1] and there is a lock-free, linearizable implementation of a queue from the same set of objects [16], neither of these implementations is strongly-linearizable. In fact, Attiya, Castañeda, and Hendler [5] proved that no lock-free, strongly-linearizable implementation of a stack or queue from such objects is possible.
The challenge when implementing a strongly-linearizable bag from these objects is for an operation that is trying to take an element from the bag to detect when the bag is empty. If elements can reside in multiple locations, it does not suffice for the operation to simply examine these locations one by one, concluding that the bag is empty if no elements are found. The problem is that new elements may have been added to locations after they were examined. Even if there was a configuration during the operation in which the bag was empty, it might not be possible to linearize the operation at this point while maintaining strong linearizability. For example, after this point, in an alternative execution, a value could be added to a location that the operation had not yet examined and the operation could return that value, instead of EMPTY.
We address this challenge in Section 4, modifying the lock-free, linearizable implementation of a queue by adding an additional readable fetch&increment object. This object is used by operations which are inserting elements into the bag to inform operations which are trying to take elements from the bag that the bag is no longer empty. We prove that the resulting algorithm is a strongly-linearizable implementation of a bag. Interestingly, although our algorithm is still a linearizable implementation of a queue, it is not a strongly-linearizable implementation of a queue.
A bag can grow arbitrarily large, so either the number of objects used to implement it or the size of the objects must be unbounded. An alternative is a $b$-bounded bag, which can simultaneously contain up to $b$ elements. Unfortunately, a $b$-bounded bag shared by more than $2b$ processes has no lock-free, strongly-linearizable implementation from interfering objects. However, we provide a lock-free, strongly-linearizable implementation of a restricted version of a $b$-bounded bag, into which only one process, the producer, can insert elements. It uses a bounded number of registers, each of bounded size, and readable, resettable test&set objects. Before presenting this implementation, we give two different implementations of a 1-bounded bag with a single producer, which introduce some of the ideas that appear in our $b$-bounded bag implementation.
In Section 5, we present a wait-free, linearizable implementation of a 1-bounded bag. Note that it provides a stronger progress guarantee. To keep the space bounded, when the producer inserts an element into the bag, it may reuse objects previously used for other elements. It has to do this carefully, to avoid reusing objects that other processes may be poised to access. This requires delicate coordination between the producer and the consumers.
In Section 6, we present a lock-free, strongly-linearizable implementation of a 1-bounded bag. It combines the mechanism for detecting an empty bag from our lock-free, strongly-linearizable implementation of a bag and the mechanism for reusing objects from our wait-free, linearizable implementation of a 1-bounded bag.
Our lock-free, strongly-linearizable implementation of a $b$-bounded bag, for any value of $b$, is presented in Section 7. To enable the producer to detect when the bag is full, we use an approach that is similar to detecting when the bag is empty, but we have to ensure that these two mechanisms do not interfere with one another. In our implementation of a 1-bounded bag in Section 6, the producer determines whether the bag is empty or full by inspecting a single object. This makes this implementation both simpler and more efficient than the special case of our implementation of a $b$-bounded bag when $b=1$.
2 Preliminaries
We consider an asynchronous system where processes communicate through operations applied to shared objects. Each type of object supports a fixed set of operations that can be applied to it. For example, a register supports read(), which returns the value of the register, and write($x$), which changes the value of the register to $x$.
Each process can be modelled as a deterministic state machine. A step by a process specifies an operation that the process applies to a shared object and the response, if any, that the operation returns. The process can then perform local computation, perhaps based on the response of the operation, to update its state. A process can crash, after which it takes no more steps.
A configuration consists of the state of every process and the value of every shared object. In an initial configuration, each process is in its initial state and each object has its initial value. An execution is an alternating sequence of configurations and steps, starting with an initial configuration. If an execution is finite, it ends with a configuration. The order in which processes take steps is determined by an adversarial scheduler.
The consensus number of an object is the largest positive integer $n$ such that there is an algorithm solving consensus among $n$ processes using only instances of this object and registers. A register has consensus number 1. Another example of an object with consensus number 1 is an ABA-detecting register [2]. It supports dWrite($x$) and dRead(). When dWrite($x$) is performed, $x$ is stored in the object. When a process $p$ performs dRead(), the value stored in the object is returned, together with an output bit, which is true if and only if process $p$ has previously performed dRead() and some dWrite() has been performed since its last dRead(). We use a restricted version of an ABA-detecting register, where dWrite() does not have an input parameter and dRead() simply returns the output bit.
A test&set object has initial value 0 and supports only one operation, t&s(). If the value of the object was 0, this operation changes its value to 1 and returns 0. If the value of the object was 1, it just returns 1. In the first case, we say that the t&s was successful and, in the second case, we say that it was unsuccessful. A resettable test&set object is like a test&set object, except that it also supports reset(), which changes the value of the object to 0. A fetch&increment object supports f&i(), which increases the value of the object by 1 and returns the previous value of the object. A compare&swap object supports the operation c&s($u,v$), which checks whether the object has value $u$ and, if so, the operation is successful and it changes the value of the object to $v$. It always returns the value of the object immediately before the operation was performed. For any type of object that does not support read, its readable version supports read in addition to its other operations.
Two operations, $op$ and $op^{\prime}$, commute if, whenever $op$ is applied to the object followed by $op^{\prime}$, the resulting value of the object is the same as when $op^{\prime}$ is applied followed by $op$. Note that the responses to $op$ and $op^{\prime}$ may be different when these operations are applied in the opposite order. The operation $op^{\prime}$ overwrites the operation $op$ if, whenever $op$ is applied to the object followed by $op^{\prime}$, the resulting value of the object is the same as when only $op^{\prime}$ is applied. Note that the response to $op^{\prime}$ may be different depending on whether $op$ was applied. An object is interfering if every two of its operations either commute or one of them overwrites the other. All interfering objects have consensus number at most 2 [12]. Registers are interfering. The test&set, resettable test&set, and fetch&increment objects and their readable versions are all interfering and have consensus number 2.
Both a stack and a queue are also examples of objects with consensus number 2 [12]. However, they are not interfering. The value of these objects is an unbounded sequence of elements, which is initially empty. A stack supports push($x$) and pop(), whereas a queue supports enqueue($x$) and dequeue(). A pop() or dequeue() that returns a value is successful. If it returns EMPTY, it is unsuccessful.
A bag is another object that is not interfering. Its value is a multiset of elements, that is initially empty. The number of elements that can be in the multiset is unbounded. The elements are taken from a set of values $V$ that does not include $\bot$. A bag supports two operations, Insert($x$) where $x\in V$ and Take. When Insert($x$) is performed, the element $x$ is added to the multiset. It does not return anything. If the multiset is not empty, a Take operation removes an arbitrary element from the multiset and returns it. In this case, we say that the operation is successful. If the multiset is empty, then Take returns EMPTY and we say that it is unsuccessful. Note that this specification is nondeterministic.
A b-bounded bag object is an object whose value is a multiset of at most $b$ elements. If the multiset contains fewer than $b$ elements when Insert($x$) is performed, the element $x$ is added to the multiset and OK is returned. If the multiset already contains $b$ elements, the value of the bag does not change and FULL is returned. For simplicity, we assume no value repetitions in the bounded bags in our proofs.
An implementation of an object $O$ shared by a set of processes $P$ from a collection of objects $C$ provides a representation of $O$ using objects in $C$ and algorithms for each process in $P$ to perform each operation supported by $O$. In an execution, an operation on an implemented object begins when a process performs the first step of its algorithm for performing this operation and is completed at the step in which the process returns from the algorithm, with the response of the operation, if any.
A bag has a straightforward implementation from a stack or a queue. The multiset is represented by the sequence, Insert($x$) is performed by applying push($x$) or enqueue($x$) and Take() is performed by applying pop() or dequeue(). Thus a bag has consensus number at most 2.
There is a simple algorithm for solving consensus among 2 processes using two single-writer registers and a bag initialized with one element: To propose the value $x$, a process writes $x$ to its single-writer register and then performs a Take() operation on the bag. If the Take() returns an element, the process returns $x$. Otherwise, it reads and returns the value that the other process wrote to its single-writer register. It follows from Borowsky, Gafni, and Afek’s Initial State Lemma [7, Lemma 3.2], that consensus among 2 processes can be solved from registers and two initially empty bags. Thus, a bag has consensus number exactly 2.
For any execution, $\alpha$, consider an ordering of every completed operation and a (possibly empty) subset of the operations that have begun, but have not completed, such that, for every completed operation, $op$, each of these operations that begins after $op$ has completed is ordered after $op$. A way to obtain such an ordering is to first assign a linearization point in $\alpha$ to each of these operations within its operation interval, i.e., no operation is linearized before it begins or after it has completed. Then specify an ordering of the operations that are linearized at the same point. Suppose there is a sequential execution in which the operations in the ordering are performed in order such that every completed operation in $\alpha$ has the same response as it does in this sequential execution. Then we say that the ordering is a linearization of $\alpha$.
An implementation of an object is linearizable [14] if each of its executions has a linearization. An implementation is strongly-linearizable [10] if there is a function $f$ that maps each execution $\alpha$ to a linearization of $\alpha$ such that, for every prefix $\alpha^{\prime}$ of $\alpha$, $f(\alpha^{\prime})$ is a prefix of $f(\alpha)$.
An implementation is wait-free if, in every execution, every operation by a process that does not crash completes within a finite number of steps by that process. An implementation is lock-free if every infinite execution contains an infinite number of completed operations.
3 Related Work
Many universal constructions (for example, from compare&swap objects and registers or from consensus objects and registers) provide strongly-linearizable, wait-free implementations of any shared object [10].
Treiber [21, 19] gave a simple lock-free, strongly-linearizable implementation of a stack from a readable compare&swap object and registers. Each push or successful pop performs exactly one successful c&s, which is where the operation is linearized. Each unsuccessful pop performs exactly one read of a compare&swap object that returns NULL, which is where the operation is linearized. Michael and Scott [18] gave a lock-free, linearizable implementation of a queue from readable compare&swap objects and registers. Assuming nodes are never reused, it is possible to make a small modification to their implementation to make it strongly-linearizable [15]. Each enqueue or successful dequeue performs exactly one successful c&s to an object other than $Tail$, which is where it is linearized. Each unsuccessful dequeue performs exactly one read of a compare&swap object that returns NULL, which is where the operation is linearized.
Even though there are wait-free, linearizable implementations of registers, max-registers, and single-writer snapshot objects from single-writer registers, Helmi, Higham, and Woelfel [11] proved that there are no lock-free, strongly-linearizable implementations. They also gave a wait-free, strongly-linearizable implementation of a bounded max-register from registers. Denysyuk and Woelfel [9] proved there are no wait-free, strongly-linearizable implementations of max-registers and single-writer snapshot objects from registers, although they gave lock-free strongly-linearizable implementations. Later, Ovens and Woelfel [20] gave lock-free, strongly-linearizable implementations of an ABA-detecting register and a single-writer snapshot object from a bounded number of bounded size registers.
Afek, Gafni, and Morrison [1] gave a wait-free, linearizable implementation of a stack from interfering objects (specifically, a fetch&add object, registers, and test&set objects). Attiya and Enea [6] showed that this implementation is not strongly-linearizable. More generally, Attiya, Castañeda, and Hendler [5] proved that there is no lock-free, strongly-linearizable implementation of a stack or a queue shared by more than 2 processes from interfering objects.
Li [16] gave a lock-free, linearizable implementation of a queue from a fetch&increment object, an infinite array of test&set objects, and an infinite array of registers. This implementation appears in Figure 1. He also gave a wait-free, linearizable implementation of a queue in which at most 2 processes may perform dequeue. It uses a fetch&increment object, an infinite array of registers, and each dequeue dynamically allocates a constant number of registers and test&set objects. David [8] gave a wait-free, linearizable implementation of a queue in which at most 1 process may enqueue. It uses registers, an infinite array of fetch&increment objects, and a two-dimensional infinite array of swap objects. It is unknown whether there is a wait-free, linearizable implementation of a queue from interfering objects in which any number of processes can enqueue and dequeue.
1Shared variables:
2 Items[1..]: an infinite array of registers, each initialized to $\bot$
3 TS[1..]: an infinite array of test&set objects, each initialized to 0
4 Max: a readable fetch&increment object, initialized to 1
5Insert(x):
6 max $\leftarrow$ Max.f&i()
7 Items[max].write(x)
8Take():
9 taken_old $\leftarrow$ 0
10 max_old $\leftarrow$ 0
[…]
Figure 1: An implementation of a queue that is not a strongly-linearizable bag
Attiya, Castañeda, and Enea [3] claimed that Li’s lock-free implementation of a queue was also a strongly-linearizable implementation of a bag. However, their claim is false.
Let $ins_{1}$ be a call of Insert(1), let $ins_{2}$ be a call of Insert(2), and let $tk$ be a call of Take(). Consider the following execution $\alpha$ of their algorithm:
  • $ins_{1}$ performs its f&i of Max on Line 6.
  • $ins_{2}$ performs its f&i of Max on Line 6.
  • tktk performs Line 9, Line 10, and does an entire iteration of the repeat loop starting on LABEL:nonSL:repeat, in which it sets max_new to 2 on LABEL:nonSL:take:read_max and sets max_old to 2 on line LABEL:nonSL:update_max_old.
  • tktk starts a second iteration of the repeat loop, in which it sets max_new to 2 on LABEL:nonSL:take:read_max and reads $\bot$ from Items[1] on LABEL:nonSL:take:read_item. However, it does not read Items[2] on LABEL:nonSL:take:read_item.
  • $ins_{1}$ writes 1 to Items[1] on Line 7 and returns.
Since $ins_{1}$ completes in $\alpha$, it must be linearized in $\alpha$.
Let $\alpha_{1}$ be the continuation of $\alpha$ in which $tk$ runs by itself to completion. It returns EMPTY as it encounters $\bot$ in every entry of Items it reads. Since $tk$ returns EMPTY in $\alpha\cdot\alpha_{1}$, it must be linearized before $ins_{1}$. Hence, $tk$ must be linearized in $\alpha$.
Let $\alpha_{2}$ be another continuation of $\alpha$, in which $ins_{2}$ writes 2 to Items[2] on Line 7 and returns, and then $tk$ runs by itself to completion. From the code, $tk$ reads 2 from Items[2] and returns 2. This implies that $ins_{2}$ must be linearized before $tk$ in $\alpha\cdot\alpha_{2}$ and, hence, in $\alpha$. Therefore $tk$ returns 2 in $\alpha\cdot\alpha_{1}$, which is a contradiction.
This shows that the algorithm in Figure 1 is not a strongly-linearizable implementation of a bag for any choice of linearization points. In particular, there is a problem with the linearization points mentioned in [4]:
  • Each Insert operation is linearized when it performs Line 7.
  • Each successful Take operation is linearized when it performs a successful test&set on LABEL:nonSL:take:t&s.
  • Each unsuccessful Take operation is linearized when it last reads Max on LABEL:nonSL:take:read_max.
The issue is that, when an unsuccessful Take operation is linearized, the bag might not be empty. Let insins be a call of Insert(1) and let $tk_{1}$ and $tk_{2}$ be calls of Take(). Consider the following execution:
  • insins performs its f&i of Max on Line 6.
  • $tk_{1}$ performs Line 9, Line 10, and does an entire iteration of the repeat loop starting on LABEL:nonSL:repeat in which it sets max_new to 1 on LABEL:nonSL:take:read_max and sets max_old to 1 on line LABEL:nonSL:update_max_old.
  • insins writes 1 to Items[1] on Line 7 and returns.
  • $tk_{1}$ starts a second iteration of the repeat loop, in which it sets max_new to 1 on LABEL:nonSL:take:read_max.
  • $tk_{2}$ performs a Take operation, returning the value 1 when it first performs LABEL:nonSL:take:t&s.
  • $tk_{1}$ completes its Take operation, returning EMPTY when it returns on LABEL:nonSL:take:return_EMPTY.
Note that, when $tk_{1}$ last performs LABEL:nonSL:take:read_max, where it is linearized, the bag contains the element 1.
A $b$-bounded bag shared by more than $2b$ processes has no lock-free, strongly-linearizable implementation from interfering objects. To see why, suppose $A$ is such an implementation for $n>2b$ processes. If a read operation is added to an interfering object, the result is still an interfering object. Hence, $A$ can be viewed as an implementation from readable interfering objects. Attiya, Castaneda, and Enea [3] defined a $b$-ordering object. A $b$-bounded bag is an example of a $b$-ordering object. Applying Lemma 5.2 from their paper shows that $b$-set agreement among $n$ processes can be solved using $A$. Since all interfering objects have consensus number at most 2, this implies that $b$-set agreement among $n>2b$ processes can be solved using objects of consensus number at most 2. Herlihy and Rajsbaum [13] proved that this is impossible.
4 A Lock-Free, Strongly-Linearizable Implementation of a Bag
The first algorithm we present is a strongly-linearizable implementation of an unbounded bag. The challenge in designing such an algorithm from interfering objects lies in identifying when the bag is empty. To address this, we use a readable fetch&increment object, Done, which an Insert operation increments as its last step, to inform Take operations that the bag is not empty. A detailed description of the algorithm follows.
The implementation uses an infinite array of registers, $Items$, in which elements that have been inserted into the bag are stored, together with an infinite array of test&set objects, $TS$. The process that performs a successful t&s() on $TS[i]$ returns the element stored in $Items[i]$ and removes it from the bag.
Each Insert($x$) operation begins by performing f&i() on a readable fetch&increment object, Allocated. This allocates the operation a new location within Items in which it writes $x$. Thus, the value of Allocated is the index of the last allocated location within Items. Finally, the operation performs f&i() on a second readable fetch&increment object, Done, to inform Take() operations about the insertion.
A Take() operation begins by reading Done and Allocated. Then it reads the allocated locations in Items. For each location, i, if Items[i] contains $x\neq\bot$, the operation performs t&s() on the corresponding test&set object, TS[i]. If this is successful, $x$ is returned. After reading all the allocated locations in Items without performing a successful t&s(), the operation rereads Done. If its value has not changed since the operation last read Done, EMPTY is returned. Otherwise, the operation repeats the entire sequence of steps. Note that it begins again starting from the first location, in case an Insert operation that was allocated an early location has recently written to that location. Pseudocode for our implementation of a bag appears in Figure 2.
1Shared variables:
2 Items[1..]: an infinite array of registers, each initialized to $\bot$
3 TS[1..]: an infinite array of test&set objects, each initialized to 0
4 Allocated: a readable fetch&increment object, initialized to 0
5 Done: a readable fetch&increment object, initialized to 0
6Insert(x):
7 m $\leftarrow$ Allocated.f&i() + 1
8 Items[m].write(x)
9 Done.f&i()
10Take():
11 repeat
12 d $\leftarrow$ Done.read()
13 m $\leftarrow$ Allocated.read()
14 for i $\leftarrow$ 1 to m do
15 x $\leftarrow$ Items[i].read()
16 if x $\neq$ $\bot$ then
[…]
Figure 2: A lock-free, strongly-linearizable bag
Strong-linearizability. Consider any execution consisting of operations on this data structure. We linearize the operations as follows:
  • A Take() operation that performs a successful t&s() on LABEL:unbounded:take:t&s is linearized at this step. It returns the element it last read on Line 15.
  • A Take() operation that reads Done on LABEL:unbounded:take:reread_Done and obtains the same value it obtained in its previous read of Done (on Line 12) is linearized at this last read. It returns EMPTY.
  • Consider an Insert($x$) operation, $ins,$ that was allocated location m and wrote $x$ to Items[m] on Line 8. If some Take() operation performs a successful t&s() on TS[m] after insins performed Line 8, but before insins performs Done.f&i() on Line 9, then insins is linearized immediately before this Take(). In this case, we say that these two operations are coupled. Otherwise, insins is linearized when it performs Done.f&i().
The linearization point of each operation occurs within its operation interval. At any configuration, $C$, the bag contains the difference between the multiset of elements inserted by Insert operations linearized before $C$ and the multiset of elements returned by successful Take operations linearized before $C$.
Each Take() operation is linearized immediately before it returns. Thus, if it is linearized at some point in an execution, it is linearized at the same point in every extension of that execution. Each Insert($x$) operation that is coupled with a Take() operation is linearized immediately before that Take() operation. This means that the inserted element is taken from the bag immediately after it is inserted. Thus, if a coupled Insert($x$) operation is linearized at some point in an execution, it is linearized at the same point in every extension of that execution. Each uncoupled Insert($x$) operation is linearized when it increments Done. This ensures that, if a Take() operation sees that the value of Done has not changed between Line 12 and LABEL:unbounded:take:reread_Done, then no uncoupled Insert has been linearized during this part of the execution. An uncoupled Insert($x$) operation is linearized immediately before it returns, so it is linearized at the same point in every extension of the execution. Hence, if the implementation is linearizable, it is strongly-linearizable.
To show that the ordering is a linearization, it remains to prove that the values returned by the Take() operations are consistent with the sequential specifications of a bag. Let $tk$ be a Take() operation that performs a successful t&s() of TS[$i$] on LABEL:unbounded:take:t&s and let $x\neq\bot$ be the element it read from Items[$i$] when it last performed Line 15. Since $tk$ read $x$ from Items[$i$], there was an Insert($x$) operation that wrote $x$ into Items[$i$] on Line 8. Note that there was exactly one Insert operation that was allocated $i$ on Line 7, by the semantics of f&i(). This Insert($x$) operation was linearized before $tk$, either because it executed Line 9 before $tk$ performed its successful t&s() on TS[$i$] or it was coupled with $tk$ and, hence, linearized immediately before $tk$. This occurrence of element $x$ remains in the bag until the linearization point of $tk$ because the element in Items[$i$] is only returned by the Take() operation that performs the successful t&s() on TS[$i$].
Now let $tk$ be a Take operation that reads Done on LABEL:unbounded:take:reread_Done and obtains the same value it obtained in its previous read of Done on Line 12. Let $C$ be the configuration immediately after $tk$ reads Done on Line 12 in its last iteration of the repeat loop and let $C^{\prime}$ be the configuration immediately before $tk$ reads Done for the last time on LABEL:unbounded:take:reread_Done. The value of Done does not change between these two configurations. Let $m$ be the value $tk$ read from Allocated on Line 13, which is the largest location in Items that had been previously allocated to an Insert operation.
Any element, $x$, that was in the bag in configuration $C$ was inserted into the bag by an uncoupled Insert($x$) operation. This operation was linearized when it performed Line 9. Suppose that it wrote $x$ into Items[$i$]. This happened prior to $C$. Note that $i\leq m$, since $m$ was obtained when $tk$ read Allocated on Line 13 after $C$ and the value of Allocated only increases. When $tk$ performed Line 15 between configurations $C$ and $C^{\prime}$, it read $x$ from Items[$i$]. Since $tk$ later reaches LABEL:unbounded:take:reread_Done, its t&s() on TS[$i$] was unsuccessful. Some other Take operation previously performed a successful t&s() on TS[$i$] and, when it did, this occurrence of element $x$ was taken from the bag. Therefore, all elements that were in the bag in configuration $C$ were taken by Take operations other than $tk$ before $C^{\prime}$.
Now consider any element that was inserted into the bag between configurations $C$ and $C^{\prime}$ by some Insert operation. Since the value of Done did not change between $C$ and $C^{\prime}$, this Insert operation did not perform Done.f&i() on Line 9. Hence, it was coupled with a Take operation and it was immediately taken from the bag after it was inserted into the bag. Therefore, all elements that were inserted into the bag between configurations $C$ and $C^{\prime}$ were taken from the bag prior to configuration $C^{\prime}$. Hence, the bag is empty in configuration $C^{\prime}$, immediately before the step at which $tk$ is linearized.
Lock-freedom. The Insert operation performs 3 steps, so it is wait-free. When a Take operation finishes an iteration of the loop and is about to begin another iteration, the value of Done it last read on LABEL:unbounded:take:reread_Done was different than what it last read on Line 12. This means that some Insert operation performed Done.f&i() on Line 9 between these two points of the execution. This step completes the Insert operation.
4.1 A linearizable, but not strongly-linearizable, queue
It is worth noting that the algorithm in Figure 2 also implements a linearizable queue. The proof of this fact is very similar to the proof used by Li [16]. However, it is not a strongly-linearizable queue, as we next prove.
Let $ins_{1}$ be a call of Insert($1$), let $ins_{2}$ be a call of Insert($2$), and let $tk$, $tk_{1}$, and $tk_{2}$ be calls of Take(). Consider the following execution $\alpha$ of the algorithm in Figure 2:
  • $ins_{1}$ performs its f&i of Allocated on Line 7 and is allocated location 1.
  • $ins_{2}$ performs its f&i of Allocated on Line 7 and is allocated location 2.
  • tktk reads 2 from Allocated on Line 13 and then reads $\bot$ from Items[1] on Line 15.
  • $ins_{1}$ and $ins_{2}$ run to completion.
Since $ins_{1}$ and $ins_{2}$ complete in $\alpha$, they must be linearized in $\alpha$.
Let $\alpha_{1}$ be the continuation of $\alpha$ in which $tk$ runs by itself to completion. It reads $2$ from Items[2] on Line 15 during its second iteration of the for loop and returns $2$. To satisfy the semantics of a queue, $ins_{2}$ must be linearized before $ins_{1}$ in $\alpha\cdot\alpha_{1}$ and, hence, in $\alpha$.
Let $\alpha_{2}$ be another continuation of $\alpha$, in which $tk_{1}$ runs by itself to completion and then $tk_{2}$ runs by itself to completion. From the code, $tk_{1}$ reads 1 from Items[1] and returns 1, and $tk_{2}$ reads 2 from Items[2] and returns 2. To satisfy the semantics of a queue, $tk_{1}$ must be linearized before $tk_{2}$ in $\alpha\cdot\alpha_{2}$ and, hence, in $\alpha$. This is a contradiction.
5 A Wait-Free, Linearizable Implementation of a 1-Bounded Bag with a Single Producer
11Shared variables:
12 Items[1..n+1]: an array of single-writer registers, all initialized to $\bot$, which can only be written to by the producer
13 TS[1..n+1]: an array of readable, resettable test&set objects, all initialized to 0, except for TS[1], which is initialized to 1
14 Allocated: a single-writer register, initialized to 1, which can only be written to by the producer
15 Hazards[1..n]: an array of single-writer registers, all initialized to $\bot$, where Hazards[i] can only be written to by $P_{i}$
16Persistent local variables of the producer:
17 used: a register, initialized to $\emptyset$, which contains a subset of $\{1,\ldots,n+1\}$
18 m: a register, initialized to 1, which is a local copy of Allocated
19Insert(x) by the producer:
[…]
21 return x
22 Hazards[i].write($\bot$)
23 return EMPTY
Figure 3: A wait-free, linearizable 1-bounded bag with one producer and $n$ consumers
Our first bounded bag algorithm from interfering objects is a wait-free, linearizable implementation of a bag that can contain at most one element. It is shared by $n$ processes, $P_{1},\ldots,P_{n}$, called consumers, that can perform Take() and a single process, called the producer, that can perform Insert($x$). To keep the space bounded, the producer reuses locations in Items to store new elements. The producer announces the most recent location it has allocated by writing it to the shared register Allocated. There is a resettable test&set object, instead of a test&set object, associated with each location in Items. An array, Hazards, of hazard pointers [17] is used to prevent multiple consumers from returning the same element, and to ensure that each element is consumed before being overwritten by a new element. Each consumer announces a location it is about to access and the producer avoids reusing the announced locations. Specifically, the producer avoids resetting the corresponding test&set objects and writing new elements to the corresponding locations in Items.
The producer maintains two persistent local variables, m, which is the last location in Items it allocated, and used, which is a set of locations it has used and will need to reset before they are reallocated. To eliminate the need for a special case to handle the first Insert call (which is the only one not preceded by a previous insertion), the initial state of the data structure is as if location 1 had been allocated to the producer, it had inserted an element into the bag in this location, and then this element had been taken by some consumer. This is simulated by setting the initial values of m, Allocated, and TS[1] to 1.
The producer begins an Insert($x$) operation by checking whether the test&set object, TS[m], in the last allocated location, m, is 0. If so, it returns FULL. Otherwise, it overwrites the element in Items[m] with $\bot$ and adds the index m to used. Afterwards, it collects the set of hazardous locations stored in Hazards. It then allocates an arbitrary location from $\{1,\ldots,n+1\}$ that is not hazardous and announces this location by writing it to Allocated. Next, it resets the test&set for each location in used that is not hazardous. Then it removes these locations from used. Finally, it writes $x$ to the newly allocated location in Items and returns OK.
To perform a Take() operation, a consumer reads Allocated. It announces the location it read in Hazards and then reads the value in this location. If it is an element $x\neq\bot$, the consumer performs t&s() on the resettable test&set object for this location. It then clears its announcement. If the t&s() was successful, it returns $x$. If either $x=\bot$ or the t&s() was unsuccessful, it returns EMPTY.
If Items[$a$] contains an element $x\neq\bot$, but this element has already been taken from the bag, it is essential that no consumer can perform a successful t&s() on the associated test&set object, TS[$a$], and take the element again. To ensure this, a consumer, $P_{i}$ writes $a$ to Hazards[$i$] before reading Items[$a$]. This prevents the producer from resetting TS[$a$] while $P_{i}$ is poised to access TS[$a$].
After it has set Items[m] to $\bot$, but before resetting any test&set objects, the producer will collect the locations that appear in Hazards and refrain from resetting the test&set objects associated with the hazardous locations it collected. Hence, it will not reset a test&set object which any consumer is poised to access.
Assuming that the universe of elements that can be inserted into the bag is bounded, the registers used in the implementations store bounded values and thus, the amount of space used by the implementation is bounded. Pseudocode for our implementation appears in Figure 3. Because the code contains no unbounded loops, the implementation is wait-free.
5.1 Linearizability
Consider any execution consisting of operations on this data structure. We linearize the operations as follows:
  • A Take() operation that performs a successful t&s() on LABEL:WF-1-bounded:take:t&s is linearized at this step. It returns the element it last read on LABEL:WF-1-bounded:take:read_item.
  • Consider a Take() operation, $tk$, that returns EMPTY. If Items[$a$] = $\bot$ or TS[$a$] = 1 when $tk$ read $a$ from Allocated on LABEL:WF-1-bounded:take:read_Allocated, then $tk$ is linearized at this step. Otherwise, we can show that some other Take() operation performed a successful t&s() on TS[$a$] after $tk$ performed LABEL:WF-1-bounded:take:read_Allocated, but before $tk$ returned. In this case, $tk$ is linearized immediately after the first such Take() operation.
  • An Insert($x$) operation that reads 0 from TS[m] (on LABEL:WF-1-bounded:insert:check_TS[m]) is linearized at this read. It returns FULL.
  • An Insert($x$) operation that writes $x$ on LABEL:WF-1-bounded:insert:write_item is linearized at this write. It returns OK.
Keeping track of hazardous locations enables objects to be safely reused so that bounded space is used. However, this makes the algorithm and its proof of correctness more intricate. We now prove that Figure 3 is a linearizable implementation of a 1-bounded bag. We begin by proving a number of facts about the algorithm to facilitate our proof.
First, we show that every location of Items contains $\bot$, except for when the producer is in a certain part of its code. When it is in that part of its code, only location Allocated contains an element.
Lemma 1
Before the producer first performs LABEL:WF-1-bounded:insert:write_item and when the producer is strictly between LABEL:WF-1-bounded:insert:clear_item and LABEL:WF-1-bounded:insert:write_item, all entries of Items are $\bot$. Otherwise, Items[Allocated] $\neq\bot$ and all other entries of Items are $\bot$.
Proof 1
All components of Items are initialized to $\bot$. The first successful Insert writes an element to Items[Allocated] on LABEL:WF-1-bounded:insert:write_item. Any subsequent successful Insert writes $\bot$ to Items[Allocated] on LABEL:WF-1-bounded:insert:clear_item. Note that the producer does not change the value of Allocated between LABEL:WF-1-bounded:insert:write_item and LABEL:WF-1-bounded:insert:clear_item, so, immediately after LABEL:WF-1-bounded:insert:clear_item, all components of Items are $\bot$. Later, on LABEL:WF-1-bounded:insert:write_item, the producer writes an element to Items[Allocated].
Next, we show that when a location in TS can be reset, the corresponding location in Items contains $\bot$.
Lemma 2
If $k$ $\in$ used, then Items[k] = $\bot$.
Proof 2
The producer adds the location m = $k$ to used on LABEL:WF-1-bounded:insert:add_to_used, after it sets Items[m] to $\bot$ on LABEL:WF-1-bounded:insert:clear_item. Thus, Items[k] = $\bot$ when $k$ is added to used. Suppose that an element is subsequently written to Items[k]. This occurs on LABEL:WF-1-bounded:insert:write_item, with m = $k$. The value of m was chosen on LABEL:WF-1-bounded:insert:choose_index to be a value that is not in hazardous. Thus, if $k$ was in used, it was removed from used on LABEL:WF-1-bounded:insert:narrow_used_down, before an element was written to Items[m].
When a Take operation returns an element, we can identify an interval immediately before it performed its successful t&s in which this test&set object was 0 and the corresponding location in Items contained the value it returned.
Lemma 3
Suppose $tk$ is a Take operation that performs a successful t&s() of TS[$a$] on LABEL:WF-1-bounded:take:t&s. Let $x$ be the element that $tk$ last read from Items[$a$] on LABEL:WF-1-bounded:take:read_item. Then TS[$a$] = 0 and Items[$a$] = $x$ between when the producer last wrote to Items[$a$] prior to the read and when $tk$ performed its successful t&s().
Proof 3
Let $D_{2}$ be the configuration right before this step, let $D_{1}$ be the configuration right before $tk$ reads Items[a] on LABEL:WF-1-bounded:take:read_item, and let $D_{0}$ be the configuration right after the producer last wrote to Items[a] on LABEL:WF-1-bounded:insert:write_item before $D_{1}$. Let $P_{i}$ be the process that performed $tk$. Prior to reading $x$ from Items[a] right after $D_{1}$, process $P_{i}$ wrote $a$ to Hazards[i] on LABEL:WF-1-bounded:take:add_hazard. It next writes to Hazards[i] on LABEL:WF-1-bounded:take:clear_hazard1, which occurs after $D_{2}$. So, between $D_{1}$ and $D_{2}$, Hazards[i] = $a$. Since $tk$ reads $x$ from Items[a] immediately after $D_{1}$ and the producer does not write to Items[a] on LABEL:WF-1-bounded:insert:write_item between $D_{0}$ and $D_{1}$, Items[a] = $x$ throughout this interval.
Right after $D_{2}$, a t&s of TS[a] succeeds, so TS[a] = 0 at $D_{2}$. Assume, for sake of contradiction, that TS[a] = 1 at some point between $D_{0}$ and $D_{2}$. Then, after this point, but before $tk$ performed its successful t&s on TS[a], the producer must have reset TS[a] on LABEL:WF-1-bounded:insert:reset_TS. By Lemma 1, Items[a] = $\bot$ when the reset was performed. Thus, between $D_{1}$ and $D_{2}$, the producer must have written $\bot$ to Items[a] on LABEL:WF-1-bounded:insert:clear_item and then performed the reset. Between these two steps, the producer collects Hazards and reads $a$ from Hazards[i]. But, by LABEL:WF-1-bounded:insert:reset_loop, the producer does not reset TS[a] on LABEL:WF-1-bounded:insert:reset_TS. This is a contradiction. Thus TS[a] = 0 between $D_{0}$ and $D_{2}$.
After writing $x$ to Items[a] just before $D_{0}$, the producer next writes to Items[a] on LABEL:WF-1-bounded:insert:clear_item. Between these two writes, the producer must read 1 from TS[a] on LABEL:WF-1-bounded:insert:check_TS[m]. But TS[a] = 0 between $D_{0}$ and $D_{2}$. Thus, Items[a] = $x$ between $D_{0}$ and $D_{2}$.
It easily follows that, when a successful t&s is performed on a location in TS, the corresponding location in Items does not contain $\bot$.
Corollary 1
When a Take operation performs a successful t&s on TS[a], Items[a] $\neq\bot$.
Now we show that when the producer writes an element into some location in Items, that value is available to be taken by a consumer.
Lemma 4
When the producer writes an element to Items[m] on LABEL:WF-1-bounded:insert:write_item, the value of TS[m] is 0.
Proof 4
Let $e$ be a step of the execution in which the producer performs LABEL:WF-1-bounded:insert:write_item. Let $a$ denote the value of m at this step.
Suppose that, prior to $e$, the producer has not performed LABEL:WF-1-bounded:insert:write_item with m = $a$. Since all locations of Items initially contain $\bot$ and elements are only written into Items on LABEL:WF-1-bounded:insert:write_item, Items[a] = $\bot$ prior to $e$. Since t&s on TS[a] is only performed on LABEL:WF-1-bounded:take:t&s after seeing that Items[a] $\neq\bot$, no t&s on TS[a] was performed prior to $e$. If $a\neq 1$, then TS[a] is initially 0, so it is 0 when $e$ is performed and we are done. If $a=1$, then $a$ was added to used when the producer first performed LABEL:WF-1-bounded:insert:add_to_used.
Otherwise, consider the last step prior to $e$ in which the producer performed LABEL:WF-1-bounded:insert:write_item with m = $a$. When the producer next performs LABEL:WF-1-bounded:insert:add_to_used, it adds $a$ to used.
So in both cases $a$ was added to used prior to $e$. By Lemma 2, Items[a] = $\bot$ while $a$ was in used. Immediately after step $e$, Items[a] $\neq\bot$, so $a$ is not in used. Prior to being removed from used on LABEL:WF-1-bounded:insert:narrow_used_down, $a$ was not in hazardous, so TS[a] was reset to 0 on LABEL:WF-1-bounded:insert:reset_TS by some step $e^{\prime}$.
By Lemma 1, Items[a] = $\bot$ at $e^{\prime}$. Between $e^{\prime}$ and $e$, there was no write to Items[a] on LABEL:WF-1-bounded:insert:write_item, so Items[a] = $\bot$. When a successful t&s on TS[a] is performed, Corollary 1 implies that Items[a] $\neq$ $\bot$. Thus, no successful t&s on TS[a] was performed between $e^{\prime}$ and $e$. This implies that TS[a] = 0 when $e$ was performed.
We show that the linearization point of a Take operation that returns EMPTY occurs within its execution interval.
Lemma 5
Let $tk$ be a Take operation that returns EMPTY. Let $e$ be the step of the execution in which $tk$ reads Allocated on LABEL:WF-1-bounded:take:read_Allocated into $a$. Suppose that Items[a] $\neq$ $\bot$ and TS[a] = 0 at $e$. Then a successful t&s of TS[a] occurs after $e$ and before $tk$ completes.
Proof 5
If $tk$ reads an element from Items[a] on LABEL:WF-1-bounded:take:read_item then, as it returns EMPTY, its t&s on LABEL:WF-1-bounded:take:t&s must fail. This implies that another Take operation has performed a successful t&s on TS[a] after $e$ and before $tk$ failed its t&s, so we are done.
Otherwise, $tk$ reads $\bot$ from Items[a]. But when $tk$ executes step $e$, Items[a] $\neq$ $\bot$. By Lemma 1, at that point the producer must be strictly between LABEL:WF-1-bounded:insert:write_item and the following execution of LABEL:WF-1-bounded:insert:clear_item. Let $e^{\prime}$ be the step in which the producer entered this interval, namely, the last step of the execution in which it writes an element to Items on LABEL:WF-1-bounded:insert:write_item before $e$. As the producer does not reach LABEL:WF-1-bounded:insert:clear_item between $e^{\prime}$ and $e$, it specifically does not reach LABEL:WF-1-bounded:insert:write_to_Allocated (in which it modifies Allocated) during that interval. Hence, it wrote to location $a$ at $e^{\prime}$. By Lemma 4, TS[a] = 0 at $e^{\prime}$. TS[a] = 0 at $e$ by assumption. It must remain 0 between $e^{\prime}$ and $e$, because if it had been set to 1, the producer would have had to reset it before $e$, but the producer does not reach LABEL:WF-1-bounded:insert:reset_TS during this interval. Since $tk$ reads $\bot$ from Items[a] after $e$, it indicates that the producer previously wrote $\bot$ to Items[a] on LABEL:WF-1-bounded:insert:clear_item. Before this write (and after $e^{\prime}$) the producer read 1 from TS[a] on LABEL:WF-1-bounded:insert:check_TS[m]. This implies that a successful t&s of TS[a] was executed after $e^{\prime}$. We have shown it could not occur between $e^{\prime}$ and $e$, so it occurred between $e$ and the execution of LABEL:WF-1-bounded:take:read_item by $tk$.
The linearization point of each operation occurs at one of its own steps, except for an unsuccessful Take operation $tk$ in case Items[a] $\neq$ $\bot$ and TS[a] = 0 when it reads Allocated. In this case, it is linearized after this read and before it completes, by Lemma 5. Thus, in all cases, the linearization point of an operation is within its operation interval.
Fix an execution consisting of operations on the bag in Figure 3. To show that the ordering of operations that we defined is a linearization, it remains to prove that the values returned by the operations are consistent with the sequential specifications of a 1-bounded bag. We do this by induction on the length of the execution. Let $S$ be the ordering of a prefix, $\sigma$, of the execution, let $e$ be the next step of the execution, and $S^{\prime}$ be the ordering of $\sigma\cdot e$. Suppose $S$ is a linearization. Let $B$ be the set of items in the bag at the end of $S$. We will prove that $S^{\prime}$ is a linearization.
It suffices to consider the steps at which operations are linearized, namely when a Take operation performs a successful t&s or reads Allocated and when an Insert operation reads 0 from TS or writes to Items.
First, suppose that $e$ is a successful t&s of TS[a] on LABEL:WF-1-bounded:take:t&s by a Take operation, $tk$. Let $x\neq\bot$ be the value $tk$ read from Items[a] when it performed LABEL:WF-1-bounded:take:read_item. If there is a pending Take() operation, $tk^{\prime}$, at the end of $\sigma$ that read $a$ from Allocated on LABEL:WF-1-bounded:take:read_Allocated when Items[a] $\neq$ $\bot$ and TS[a] = 0, then $S^{\prime}$ is $S$ followed by $tk$ with return value $x$ and then $tk^{\prime}$ with return value EMPTY. Otherwise, $S^{\prime}$ is $S$ followed by $tk$ with return value $x$. In both cases, to show that $S^{\prime}$ is a linearization, we must show that $B=\{x\}$.
Let $D_{1}$ be the configuration right before this read and let $D_{2}$ be the configuration at the end of $\sigma$. Let $D_{0}$ be the configuration right after the last write to Items[a] on LABEL:WF-1-bounded:insert:write_item that was performed by the producer before $D_{1}$ and let insins be the Insert operation that performed this write. By Lemma 3, TS[a] = 0 and Items[a] = $x$ between $D_{0}$ and $D_{2}$. Therefore, by Lemma 1, no write to any cell of Items on LABEL:WF-1-bounded:insert:write_item is performed between $D_{0}$ and $D_{2}$, which means insins performed the last write of an element to Items in $\sigma$.
Since insins performed LABEL:WF-1-bounded:insert:write_item right before $D_{0}$, it was completed in $\sigma$. Hence insins appears in $S$. Immediately after it appears, the bag contains only $x$. Between $D_{0}$ and $D_{2}$, TS[a] = 0, so no successful t&s is performed on TS[a]. Also, Items[a] $\neq\bot$, so, by Lemma 1, Items[a’] $\bot$ for all $a^{\prime}\neq a$. By Corollary 1, Items[a’] $\neq\bot$ when a successful t&s is performed on TS[a’]. Thus, no successful t&s is performed on any location of TS between $D_{0}$ and $D_{2}$. Hence, no Take operation that removes $x$ is linearized between $D_{0}$ and $D_{2}$. Since insins is linearized right before $D_{0}$, there are no successful Take operations after insins in $S$. Therefore, at the end of $S$, the bag still contains $x$. Since $S$ is a linearization, it follows that $B=\{x\}$.
Next, suppose that $e$ is a read of Allocated on LABEL:WF-1-bounded:take:read_Allocated by a Take operation $tk$ into $a$, when Items[a] = $\bot$ or TS[a] = 1. Then $S^{\prime}$ is $S$ followed by $tk$ with return value EMPTY. To show that $S^{\prime}$ is a linearization, we must show that $B$ is empty. $B$ is not empty if and only if there is a successful Insert operation in $S$ after which there is no successful Take operation. This correlates to $\sigma$ in which the producer wrote to Items on LABEL:WF-1-bounded:insert:write_item and there is no following successful t&s on TS. We will show that this is not the case in $S$. Assume, for sake of contradiction, that $\sigma$ contains a write of an element to Items by the producer on LABEL:WF-1-bounded:insert:write_item, but no successful t&s on TS after the last such write. Let $e$ be the step in which the producer executes this last write. By Lemma 4, TS[Allocated] = 0 at $e$. Since no successful t&s follows $e$ in $\sigma$, the producer may not pass LABEL:WF-1-bounded:insert:check_TS[m]. In particular, it may not reach LABEL:WF-1-bounded:insert:write_to_Allocated and thus Allocated remains unchanged. It also does not reach LABEL:WF-1-bounded:insert:clear_item, hence (and since Allocated remains unchanged), Items[Allocated] remains non-$\bot$ until the end of $\sigma$. The fact that Allocated remains unchanged also implies that TS[Allocated] remains 0 until the end of $\sigma$. But these contradict the fact that in the end of $\sigma$, Items[Allocated] = $\bot$ or TS[Allocated] = 1. Thus, $B$ is empty.
Now, suppose $e$ is a read of 0 from TS[m] by an Insert(x) operation insins on LABEL:WF-1-bounded:insert:check_TS[m]. Then $S^{\prime}$ is $S$ followed by insins with return value FULL. To show that $S^{\prime}$ is a linearization, we must show that $B$ is not empty.
Only successful insertions can modify m and Allocated or reset components of TS. Initially, m = Allocated = 1, TS[1] = 1, and all other cells of TS are initially 0. Hence, for insins to read 0 from TS[m] on LABEL:WF-1-bounded:insert:check_TS[m], a previous successful Insert operation must have been performed before insins, changing the value of m and Allocated or resetting TS[1].
Let $ins^{*}$ be the last successful Insert operation the producer performed before insins. Since $ins^{*}$ has completed, it is linearized before insins. Since $ins^{*}$ is linearized when it performs LABEL:WF-1-bounded:insert:write_item, the producer does not perform LABEL:WF-1-bounded:insert:clear_item through LABEL:WF-1-bounded:insert:write_item strictly between the linearization points of $ins^{*}$ and insins. Note that m is only changed on LABEL:WF-1-bounded:insert:choose_index and Allocated is only changed on LABEL:WF-1-bounded:insert:write_to_Allocated, so their values do not change during this interval. There are also no resets performed during this interval, since these only occur on LABEL:WF-1-bounded:insert:reset_TS. Moreover, by Lemma 1, throughout this interval, Items[Allocated] $\neq\bot$ and all other locations of Items contain $\bot$. Corollary 1 implies that the only successful t&s performed during this interval is on TS[Allocated]. However, if there is such a successful t&s, it is not reset before insins reads TS[m], so it would read 1, rather than 0. Hence, no successful Take operation is linearized between $ins^{*}$ and insins, and so $B$ contains the item that $ins^{*}$ inserted.
Finally, suppose $e$ is a write to Items by an Insert(x) operation insins on LABEL:WF-1-bounded:insert:write_item. Then $S^{\prime}$ is $S$ followed by insins with return value OK. To show that $S^{\prime}$ is a linearization, we must show that $B$ is empty. This is clearly true if there are no Insert operations in $S$. Otherwise, let $ins^{*}$ be the last Insert operation in $S$. Let $e^{*}$ be the step of $\sigma$ in which $ins^{*}$ wrote to Items[$a^{*}$] on LABEL:WF-1-bounded:insert:write_item. By Lemma 4, TS[$a^{*}$] = 0 when $e^{*}$ occurs. Since $S$ is a linearization, right after $ins^{*}$, the bag contains one item. Only successful Insert operations modify m (on LABEL:WF-1-bounded:insert:choose_index), so m = $a^{*}$ when insins read 1 from TS[m] on LABEL:WF-1-bounded:insert:check_TS[m]. Thus, a successful t&s on TS[$a^{*}$] must have been performed between $e^{*}$ and this step. The Take operation that performed it is linearized at this t&s. It appears after $ins^{*}$ in $S$, so the bag is empty at the end of $S$.
5.2 Not Strongly-Linearizable
We show that Figure 3 is not a strongly-linearizable implementation of a 1-bounded bag with a single producer.
Let $ins_{i}$ be a call of Insert($i$) and $tk_{i}$ a call of Take(), for $1\leq i\leq 3$. Consider the following execution $\alpha$ of the algorithm in Figure 3:
  • $ins_{1}$ starts running, and it is allocated location 1 on LABEL:WF-1-bounded:insert:choose_index. It then runs to completion.
  • $tk_{1}$ runs and returns 1.
  • $tk_{2}$ reads 1 from Allocated on LABEL:WF-1-bounded:take:read_Allocated.
  • $ins_{2}$ starts running. It writes $\bot$ to Items[1] on LABEL:WF-1-bounded:insert:clear_item. It is allocated location 2 on LABEL:WF-1-bounded:insert:choose_index. It then runs to completion.
Let $\alpha_{1}$ be the continuation of $\alpha$ in which $tk_{2}$ runs by itself to completion. It returns EMPTY as it encounters $\bot$ in every entry of Items it reads including Items[1]. Since $tk_{2}$ returns EMPTY in $\alpha\cdot\alpha_{1}$, it must be linearized before $ins_{2}$. Hence, $tk_{2}$ must be linearized in $\alpha$, with return value EMPTY.
Let $\alpha_{2}$ be another continuation of $\alpha$, in which the following occur:
  • $tk_{3}$ runs and returns 2.
  • $ins_{3}$ starts running. It is allocated location 1 on LABEL:WF-1-bounded:insert:choose_index. It continues to run, writes 3 to Items[1] and returns.
  • $tk_{2}$ runs to completion.
From the code, $tk_{2}$ reads 3 from Items[1] and returns 3. Hence, $tk_{2}$ must be linearized after $ins_{3}$ in $\alpha\cdot\alpha_{2}$. This contradicts strong-linearizability.
6 A Lock-Free, Strongly-Linearizable Implementation of a 1-Bounded Bag with a Single Producer
17Shared variables:
18 Items[1..n+1]: an array of single-writer registers, all initialized to $\bot$, which can only be written to by the producer
19 TS[1..n+1]: an array of readable, resettable test&set objects, all initialized to 0, except for TS[1], which is initialized to 1
20 Allocated: a single-writer register, initialized to 1, which can only be written to by the producer
21 Hazards[1..n]: an array of single-writer registers, all initialized to $\bot$, where Hazards[i] can only be written to by $P_{i}$
22 Done: an ABA-detecting register
23Persistent local variables of the producer:
24 used: a register, initialized to $\emptyset$, which contains a subset of $\{1,\ldots,n+1\}$
25 m: a register, initialized to 1, which is a local copy of Allocated
26Insert(x) by the producer:
[…]
Figure 4: A lock-free, strongly-linearizable 1-bounded bag with one producer and $n$ consumers
In this section, we present a lock-free, strongly-linearizable implementation of a bag that can contain at most one element. It is shared by $n$ processes, $P_{1},\ldots,P_{n}$, called consumers, that can perform Take() and a single process, called the producer, that can perform Insert($x$). It combines ideas from our lock-free, strongly-linearizable implementation of an unbounded bag in Section 4 and our wait-free, linearizable implementation of a 1-bounded bag in Section 5. To keep the space bounded, we use an ABA-detecting register for Done, instead of a readable fetch&increment object. An ABA-detecting register has a lock-free, strongly-linearizable implementation from registers using bounded space [20].
An Insert($x$) operation by the producer is the same as in the wait-free linearizable implementation presented in the previous section, except the producer writes to the ABA-detecting register Done before it returns OK.
To perform a Take() operation, a consumer, $P_{i}$, reads Done and Allocated. In Hazards[i], it announces the location, a, that it read from Allocated. Then $P_{i}$ reads Items[a] and, if it contains an element $x\neq\bot$, then $P_{i}$ performs t&s() on TS[a]. Next $P_{i}$ clears its announcement. If the t&s() was successful, $P_{i}$ returns $x$. If either $x=\bot$ or the t&s() was unsuccessful, $P_{i}$ rereads Done and, if its value has not changed since $P_{i}$’s previous read, $P_{i}$ returns EMPTY. Otherwise, $P_{i}$ repeats the entire sequence of steps. Pseudocode for our implementation appears in Figure 4.
6.1 Strong-Linearizability
Consider any execution consisting of operations on this data structure. We linearize the operations as follows:
  • A Take() operation that performs a successful t&s() on LABEL:1-bounded:take:t&s is linearized when it performs this t&s(). It returns the value it read in its last execution of LABEL:1-bounded:take:read_item.
  • A Take() operation that gets false from Done.dRead() on LABEL:1-bounded:take:reread_done is linearized at this step. It returns EMPTY.
  • An Insert($x$) operation that reads 0 from TS[m] (on LABEL:1-bounded:insert:check_TS[m]) is linearized at this read. It returns FULL.
  • Consider an Insert($x$) operation, insins, that allocated location a to m on LABEL:1-bounded:insert:choose_index and wrote $x$ to Items[a] on LABEL:1-bounded:insert:write_item. If some Take() operation performs a successful t&s() on TS[a] after insins performed LABEL:1-bounded:insert:write_item, but before insins performs Done.dRead() on LABEL:1-bounded:insert:write_to_done, then insins is linearized immediately before the Take() operation that performed the first such t&s(). In this case, insins returns OK and we say that these Insert($x$) and Take() operations are coupled.
  • An Insert($x$) operation that performs Done.dWrite() on LABEL:1-bounded:insert:write_to_done, but has not already been linearized as a coupled operation, is linearized at this step. In this case, it returns OK and we say that it is uncoupled.
We now prove that the algorithm in Figure 4 is a strongly-linearizable implementation of a 1-bounded bag with a single producer.
The linearization point defined above for each operation, occurs at one of its own steps, except for coupled Insert operations, for which the linearization point is between its last two steps. Thus, in all cases, the linearization point of an operation is within its operation interval.
Each Take() operation is linearized immediately before it returns on LABEL:1-bounded:take:reread_done or when it performs a successful t&s on LABEL:1-bounded:take:t&s. Thus, if it is linearized at some point in an execution, it is linearized at the same point in every extension of that execution. Each Insert($x$) operation that is coupled with a Take() operation is linearized immediately before that Take() operation. Thus, if a coupled Insert($x$) operation is linearized at some point in an execution, it is linearized at the same point in every extension of that execution. Each uncoupled Insert($x$) operation is linearized immediately before it returns, so it is linearized at the same point in every extension of the execution. Hence, if the implementation is linearizable, it is strongly-linearizable.
We now prove a number of useful facts about the implementation. First, we show that every location of Items contains $\bot$, except for when the producer is in a certain part of its code. When it is in that part of its code, only location Allocated contains an element.
Lemma 6
Before the producer first performs LABEL:1-bounded:insert:write_item and when the producer is strictly between LABEL:1-bounded:insert:clear_item and LABEL:1-bounded:insert:write_item, all entries of Items are $\bot$. Otherwise, Items[Allocated] $\neq\bot$ and all other entries of Items are $\bot$.
Proof 6
All components of Items are initialized to $\bot$. The first successful Insert writes an element to Items[Allocated] on LABEL:1-bounded:insert:write_item. Any subsequent successful Insert writes $\bot$ to Items[Allocated] on LABEL:1-bounded:insert:clear_item. Note that the producer does not change the value of Allocated between LABEL:1-bounded:insert:write_item and LABEL:1-bounded:insert:clear_item, so, immediately after LABEL:1-bounded:insert:clear_item, all components of Items are $\bot$. Later, on LABEL:1-bounded:insert:write_item, the producer writes an element to Items[Allocated].
Next, we show that when a location in TS can be reset, the corresponding location in Items contains $\bot$.
Lemma 7
If $k$ $\in$ used, then Items[k] = $\bot$.
Proof 7
The producer adds the location m = $k$ to used on LABEL:1-bounded:insert:add_to_used, after it sets Items[m] to $\bot$ on LABEL:1-bounded:insert:clear_item. Thus, Items[k] = $\bot$ when $k$ is added to used. Suppose that an element is subsequently written to Items[k]. This occurs on LABEL:1-bounded:insert:write_item, with m = $k$. The value of m was chosen on LABEL:1-bounded:insert:choose_index to be a value that is not in hazardous. Thus, if $k$ was in used, it was removed from used on LABEL:1-bounded:insert:narrow_used_down, before an element was written to Items[m].
When a Take operation returns an element, we can identify an interval immediately before it performed its successful t&s in which this test&set object contained 0 and the corresponding location in Items contained the value that Take returned.
Lemma 8
Suppose $tk$ is a Take operation that performs a successful t&s of TS[$a$] on LABEL:1-bounded:take:t&s. Let $x$ be the element that $tk$ last read from Items[$a$] on LABEL:1-bounded:take:read_item. Then TS[$a$] = 0 and Items[$a$] = $x$ between when the producer last wrote to Items[$a$] prior to the read and when $tk$ performed its successful t&s.
Proof 8
Let $D_{2}$ be the configuration right before $tk$ performs the successful t&s of TS[a], let $D_{1}$ be the configuration right before $tk$ reads Items[a] on LABEL:1-bounded:take:read_item for the last time, and let $D_{0}$ be the configuration right after the producer last wrote to Items[a] on LABEL:1-bounded:insert:write_item before $D_{1}$. Let $P_{i}$ be the process that performed $tk$. Prior to reading $x$ from Items[a] right after $D_{1}$, process $P_{i}$ wrote $a$ to Hazards[i] on LABEL:1-bounded:take:add_hazard. It next writes to Hazards[i] on LABEL:1-bounded:take:clear_hazard1, which occurs after $D_{2}$. So, between $D_{1}$ and $D_{2}$, Hazards[i] = $a$. Since $tk$ reads $x$ from Items[a] immediately after $D_{1}$ and the producer does not write to Items[a] on LABEL:1-bounded:insert:write_item between $D_{0}$ and $D_{1}$, Items[a] = $x$ throughout this interval.
Right after $D_{2}$, a t&s of TS[a] succeeds, so TS[a] = 0 at $D_{2}$. Assume, for sake of contradiction, that TS[a] = 1 at some point between $D_{0}$ and $D_{2}$. Then, after this point, but before $tk$ performed its successful t&s on TS[a], the producer must have reset TS[a] on LABEL:1-bounded:insert:reset_TS. By Lemma 6, Items[a] = $\bot$ when the reset was performed. Thus, between $D_{1}$ and $D_{2}$, the producer must have written $\bot$ to Items[a] on LABEL:1-bounded:insert:clear_item and then performed the reset. Between these two steps, the producer collects Hazards and reads $a$ from Hazards[i]. But, by LABEL:1-bounded:insert:reset_loop, the producer does not reset TS[a] on LABEL:1-bounded:insert:reset_TS. This is a contradiction. Thus TS[a] = 0 between $D_{0}$ and $D_{2}$.
After writing $x$ to Items[a] just before $D_{0}$, the producer next writes to Items[a] on LABEL:1-bounded:insert:clear_item. Between these two writes, the producer must read 1 from TS[a] on LABEL:1-bounded:insert:check_TS[m]. But TS[a] = 0 between $D_{0}$ and $D_{2}$. Thus, Items[a] = $x$ between $D_{0}$ and $D_{2}$.
It easily follows that, when a successful t&s is performed on a location in TS, the corresponding location in Items does not contain $\bot$.
Corollary 2
When a Take operation performs a successful t&s on TS[a], Items[a] $\neq\bot$.
Now we show that when the producer writes an element into some location in Items, that value is available to be taken by a taker.
Lemma 9
When the producer writes an element to Items[m] on LABEL:1-bounded:insert:write_item, the value of TS[m] is 0.
Proof 9
Let $e$ be a step of the execution in which the producer performs LABEL:1-bounded:insert:write_item. Let $a$ denote the value of m at this step.
Suppose that, prior to $e$, the producer has not performed LABEL:1-bounded:insert:write_item with m = $a$. Since all locations of Items initially contain $\bot$ and elements are only written into Items on LABEL:1-bounded:insert:write_item, Items[a] = $\bot$ prior to $e$. Since t&s on TS[a] is only performed on LABEL:1-bounded:take:t&s after seeing that Items[a] $\neq\bot$, no t&s on TS[a] was performed prior to $e$. If $a\neq 1$, then TS[a] is initially 0, so it is 0 when $e$ is performed and we are done. If $a=1$, then $a$ was added to used when the producer first performed LABEL:1-bounded:insert:add_to_used.
Otherwise, consider the last step prior to $e$ in which the producer performed LABEL:1-bounded:insert:write_item with m = $a$. When the producer next performs LABEL:1-bounded:insert:add_to_used, it adds $a$ to used.
By Lemma 7, Items[a] = $\bot$ while $a$ was in used. Immediately after step $e$, Items[a] $\neq\bot$, so $a$ is not in used. Prior to being removed from used on LABEL:1-bounded:insert:narrow_used_down, $a$ was not in hazardous, so TS[a] was reset to 0 on LABEL:1-bounded:insert:reset_TS by some step $e^{\prime}$.
By Lemma 6, Items[a] = $\bot$ at $e^{\prime}$. Between $e^{\prime}$ and $e$, there was no write to Items[a] on LABEL:1-bounded:insert:write_item, so Items[a] = $\bot$. When a successful t&s on TS[a] is performed, Corollary 2 implies that Items[a] $\neq$ $\bot$. Thus, no successful t&s on TS[a] was performed between $e^{\prime}$ and $e$. This implies that TS[a] = 0 when $e$ was performed.
Fix an execution consisting of operations on the bag in Figure 4. To show that the ordering of operations that we defined is a linearization, it remains to prove that the values returned by the operations are consistent with the sequential specifications of a 1-bounded bag. We do this by induction on the length of the execution. Let $S$ be the ordering of a prefix, $\sigma$, of the execution, let $e$ be the next step of the execution, and $S^{\prime}$ be the ordering of $\sigma\cdot e$. Suppose $S$ is a linearization. Let $B$ be the set of items in the bag at the end of $S$. We will prove that $S^{\prime}$ is a linearization.
It suffices to consider the steps at which operations are linearized, namely when a Take operation performs a successful t&s or gets false from a dRead of Done and when an Insert operation reads 0 from TS or writes to Done.
First, suppose that $e$ is a successful t&s of TS[a] on LABEL:1-bounded:take:t&s by a Take operation, $tk$. Let $x\neq\bot$ be the value $tk$ read from Items[a] when it last performed LABEL:1-bounded:take:read_item during $\sigma$. Let $D_{1}$ be the configuration right before this read and let $D_{2}$ be the configuration at the end of $\sigma$. Let $D_{0}$ be the configuration right after the last write to Items[a] on LABEL:1-bounded:insert:write_item that was performed by the producer before $D_{1}$ and let insins be the Insert operation that performed this write. By Lemma 8, TS[a] = 0 and Items[a] = $x$ between $D_{0}$ and $D_{2}$. Therefore, by Lemma 6, no write to any cell of Items on LABEL:1-bounded:insert:write_item is performed between $D_{0}$ and $D_{2}$, which means insins performed the last write of an element to Items in $\sigma$.
If there is a pending Insert(x’) operation, $ins^{\prime}$, at the end of $\sigma$ that wrote to Items[m] on LABEL:1-bounded:insert:write_item (but has not incremented Done on LABEL:1-bounded:insert:write_to_done), then $ins^{\prime}=ins$. In this case, $S^{\prime}$ is $S$ followed by insins with return value OK and then $tk$ with return value $x$. To show that $S^{\prime}$ is a linearization, we must show that $B=\emptyset$. If there are no Insert operations in $S$, then the bag is empty at the end of $S$. Otherwise, let $ins^{*}$ be the last Insert operation in $S$. Let $e^{*}$ be the step of $\sigma$ in which $ins^{*}$ wrote to Items[$a^{*}$] on LABEL:1-bounded:insert:write_item. By Lemma 9, TS[$a^{*}$] = 0 when $e^{*}$ occurs. Since $S$ is a linearization, right after $ins^{*}$, the bag contains one item. Only successful Insert operations modify m (on LABEL:1-bounded:insert:choose_index), so m = $a^{*}$ when insins read 1 from TS[m] on LABEL:1-bounded:insert:check_TS[m]. Thus, a successful t&s on TS[$a^{*}$] must have been performed between $e^{*}$ and this step. The Take operation that performed it is linearized at this t&s. It appears after $ins^{*}$ in $S$, so the bag is empty at the end of $S$.
If there is no Insert operation that performed LABEL:1-bounded:insert:write_item in $\sigma$, but not LABEL:1-bounded:insert:write_to_done, then $S^{\prime}$ is $S$ followed by $tk$ with return value $x$. To show that $S^{\prime}$ is a linearization, we must show that $B=\{x\}$. Since insins performed LABEL:1-bounded:insert:write_item right before $D_{0}$, it also performed LABEL:1-bounded:insert:write_to_done and, thus, was completed in $\sigma$. Hence insins appears in $S$. Immediately after it appears, the bag contains only $x$. Between $D_{0}$ and $D_{2}$, TS[a] = 0, so no successful t&s is performed on TS[a]. Also, Items[a] $\neq\bot$, so, by Lemma 6, Items[a’] $\bot$ for all $a^{\prime}\neq a$. By Corollary 2, Items[a’] $\neq\bot$ when a successful t&s is performed on TS[a’]. Thus, no successful t&s is performed on any location of TS between $D_{0}$ and $D_{2}$. Hence, no Take operation that removes $x$ is linearized between $D_{0}$ and $D_{2}$. Since insins is linearized after $D_{0}$, there are no successful Take operations after insins in $S$. Therefore, at the end of $S$, the bag still contains $x$. Since $S$ is a linearization, it follows that $B=\{x\}$.
Next, suppose that $e$ is an execution of Done.dRead() on LABEL:1-bounded:take:reread_done by a Take operation $tk$, which returns false. Then $S^{\prime}$ is $S$ followed by $tk$ with return value EMPTY. To show that $S^{\prime}$ is a linearization, we must show that $B$ is empty.
Let $C$ be the configuration at the end of $\sigma$, let $C^{*}$ be the configuration immediately after $tk$ last read Done on either LABEL:1-bounded:take:read_done or LABEL:1-bounded:take:reread_done, let $S^{*}$ be the linearization of the prefix of $\sigma$ ending with $C^{*}$, and let $B^{*}$ be the set of items in the bag at the end of $S^{*}$. Since $tk$ returns on LABEL:1-bounded:take:reread_done, the producer does not write to Done between these two configurations.
Suppose there was a successful Insert operation that is in $S$, but is not in $S^{*}$. Then it was linearized between configurations $C^{*}$ and $C$. Since the producer did not write to Done on LABEL:1-bounded:insert:write_to_done between $C^{*}$ and $C$, this Insert operation was coupled with a Take operation. Hence it was immediately taken from the bag after it was inserted into the bag. Therefore, if $B^{*}$ is empty, then so is $B$.
So, suppose $B^{*}$ is not empty. Then it contains exactly one element, since $S$ is a linearization. Items inserted by coupled Insert operations are immediately taken from the bag by a Take operation linearized at the same step, so they are not in the bag in the end of the linearization of any prefix of the execution. Hence they are not in $B^{*}$. Therefore, the item, $x$, in $B^{*}$ was inserted into the bag by an uncoupled Insert operation, insins. This operation wrote $x$ to Items in some location, $a$, on LABEL:1-bounded:insert:write_item. When that occurred, TS[a] = 0 by Lemma 9. Since insins is uncoupled, no taker performs a successful t&s on TS[a] between this write and when insins writes to Done on LABEL:1-bounded:insert:write_to_done. Hence TS[a] = 0 at the linearization point of insins. No successful Insert follows insins in $S^{*}$, since $S^{*}$ is a linearization and the bag is 1-bounded. No successful Take operation follows insins in $S^{*}$, since it would remove $x$ from the bag. Thus, no successful t&s is performed between the linearization point of insins and $C^{*}$ and, so, TS[a] = 0 throughout this interval.
If $tk$ read a location $a^{\prime}\neq a$ from Allocated on LABEL:1-bounded:take:read_Allocated after $C^{*}$, then the producer wrote $a^{\prime}$ to Allocated on LABEL:1-bounded:insert:write_to_Allocated after insins was completed, which implies that the producer passed LABEL:1-bounded:insert:check_TS[m]. If $tk$ read $a$ from Allocated on LABEL:1-bounded:take:read_Allocated after $C^{*}$ and then read $\bot$ from Items[a] on LABEL:1-bounded:take:read_item, then the producer wrote $\bot$ to Items[a] on LABEL:1-bounded:insert:clear_item after insins was completed, which implies that the producer passed LABEL:1-bounded:insert:check_TS[m]. So in both cases, the producer read 1 from TS[a] on LABEL:1-bounded:insert:check_TS[m] after insins was completed. Since TS[a] = 0 between the end of insins and $C^{*}$, some Take operation performed a successful t&s on TS[a] between $C^{*}$ and $C$. The first Take to perform such a t&s follows insins in $S$ and takes $x$ from the bag, so $x\not\in B$.
Otherwise, after $C^{*}$ tktk read $a$ from Allocated and then read an element from Items[a]. In this case, $tk$ must have performed an unsuccessful t&s of TS[a] on LABEL:1-bounded:take:t&s, so another Take operation performed a successful t&s of TS[a] after $C^{*}$ but before this step. The first Take to perform such a t&s follows insins in $S$ and takes $x$ from the bag, so $x\not\in B$.
Now, suppose $e$ is a read of 0 from TS[m] by an Insert($x$) operation insins on LABEL:1-bounded:insert:check_TS[m]. Then $S^{\prime}$ is $S$ followed by insins with return value FULL. To show that $S^{\prime}$ is a linearization, we must show that $B$ is not empty.
Only successful insertions can modify m and Allocated or reset components of TS. Initially, m = Allocated = 1, TS[1] = 1, and all other cells of TS are initially 0. Hence, for insins to read 0 from TS[m] on LABEL:1-bounded:insert:check_TS[m], a previous successful Insert operation must have been performed before insins, changing the value of m and Allocated or resetting TS[1].
Let $ins^{*}$ be the last successful Insert operation the producer performed before insins. Since $ins^{*}$ has completed, it is linearized before insins. Since $ins^{*}$ is linearized after it performs LABEL:1-bounded:insert:write_item, the producer does not perform LABEL:1-bounded:insert:clear_item through LABEL:1-bounded:insert:write_item between the linearization points of $ins^{*}$ and insins. Note that m is only changed on LABEL:1-bounded:insert:choose_index and Allocated is only changed on LABEL:1-bounded:insert:write_to_Allocated, so their values do not change during this interval. There are also no resets performed during this interval, since these only occur on LABEL:1-bounded:insert:reset_TS. Moreover, by Lemma 6, throughout this interval, Items[Allocated] $\neq\bot$ and all other locations of Items contain $\bot$. Corollary 2 implies that the only successful t&s performed during this interval is on TS[Allocated]. However, if there is such a successful t&s, it is not reset before insins reads TS[m], so it would read 1, rather than 0. Hence, no successful Take operation is linearized between $ins^{*}$ and insins, and so $B$ contains the item that $ins^{*}$ inserted.
Finally, suppose $e$ is a write to Done by an uncoupled Insert($x$) operation insins on LABEL:1-bounded:insert:write_to_done. Then $S^{\prime}$ is $S$ followed by insins with return value OK. To show that $S^{\prime}$ is a linearization, we must show that $B$ is empty. This is clearly true if there are no Insert operations in $S$. Otherwise, let $ins^{*}$ be the last Insert operation in $S$. Let $e^{*}$ be the step of $\sigma$ in which $ins^{*}$ wrote to Items[$a^{*}$] on LABEL:1-bounded:insert:write_item. By Lemma 9, TS[$a^{*}$] = 0 when $e^{*}$ occurs. Since $S$ is a linearization, right after $ins^{*}$, the bag contains one item. Only successful Insert operations modify m (on LABEL:1-bounded:insert:choose_index), so m = $a^{*}$ when insins read 1 from TS[m] on LABEL:1-bounded:insert:check_TS[m]. Thus, a successful t&s on TS[$a^{*}$] must have been performed between $e^{*}$ and this step. The Take operation that performed it is linearized at this t&s. It appears after $ins^{*}$ in $S$, so the bag is empty at the end of $S$.
6.2 Lock-Freedom
The Insert operation has no unbounded loops, so it is wait-free. Suppose a Take operation, $tk$, finishes an iteration of the loop and is about to begin another iteration, because its last read of Done on LABEL:1-bounded:take:reread_done returned true. Then some Insert operation wrote to Done on LABEL:1-bounded:insert:write_to_done since when $tk$ previously read Done on LABEL:1-bounded:take:read_done or LABEL:1-bounded:take:reread_done. This write step completes the Insert operation.
7 A Lock-Free Strongly-Linearizable Implementation of a $b$-Bounded Bag with a Single Producer
Our final algorithm is a lock-free, strongly-linearizable implementation of a bag that can contain at most $b$ elements. It is shared by $n$ processes, $P_{1},\ldots,P_{n}$, called consumers, that can perform Take() and a single process, called the producer, that can perform Insert($x$). It extends our lock-free, strongly-linearizable implementation of a 1-bounded bag in Section 6. In this algorithm, Allocated is a shared register that stores a subset of at most $b$ locations, and alloc is a local variable the producer uses to update the contents of Allocated. The challenge in designing a strongly-linearizable bag algorithm from interfering objects was for an operation that is trying to take an element from the bag to detect when the bag is empty (and then return EMPTY). This is addressed using an ABA-detecting register, InsertDone, to which each Insert($x$) operation writes after writing $x$ to Items. When designing a strongly-linearizable bounded bag algorithm from interfering objects, an operation that is trying to insert an element into the bag faces a symmetrical challenge, as it needs to detect when the bag is full (and then return FULL). To address this, we use another ABA-detecting register, TakeDone, to which each Take() operation writes after performing a successful t&s(). Unsuccessful Take operations also write to TakeDone before returning, to help them be linearized before the unsuccessful Take.
23Shared variables:
24 Items[1..n+b]: an array of single-writer registers, each initialized to \ $\bot$, which can only be written to by the producer
25 TS[1..n+b]: an array of readable, resettable test&set objects, each initialized to 0
26 Allocated: a single-writer register, initialized to $\emptyset$, which contains a subset of $\{1,\ldots,n+b\}$ of size between 0 and $b$ and can only be written to by the producer
27 Hazards[1..n]: an array of single-writer registers, each initialized to \ $\bot$, where Hazards[i] can only be written to by $P_{i}$
28 InsertDone, TakeDone: ABA-detecting registers
29Persistent local variables of the producer:
30 used: a register, initialized to $\emptyset$, which contains a subset of $\{1,\ldots,n+b\}$
31 alloc: a register, initialized to $\emptyset$, which is used to update Allocated
32Insert(x) by the producer:
33 TakeDone.dRead()
34 repeat
35 for all m $\in$ alloc do
[…]
37 return EMPTY
Figure 5: A lock-free, strongly-linearizable $b$-bounded bag with one producer and $n$ consumers
The producer begins an Insert($x$) operation by reading TakeDone. Then it looks at each of the locations in TS that have been allocated. For each location $m$ that contains 1, it clears the element in that location (setting Items[m] to $\bot$), removes m from alloc, and adds m to used. If alloc contains less than $b$ locations, it adds $x$ to the bag as follows. It first collects the non-$\bot$ values from the Hazards array into the set hazardous. Afterwards, it allocates an arbitrary location, m, from $\{1,\ldots,n+b\}$, excluding those in alloc and in hazardous, adds this location to alloc, and announces it by copying alloc into Allocated. Next, for each location in used that is not in hazardous, it resets the associated test&set object and removes the location from used. Then it writes $x$ to the allocated object, Items[m]. Finally, it writes to the ABA-detecting register InsertDone and returns OK. If alloc contained $b$ locations, it rereads TakeDone and, if its value has not changed since its previous read, it returns FULL; otherwise, it repeats the entire sequence of steps.
To perform a Take() operation, a process reads InsertDone and Allocated. For each location $a$ in the set of locations obtained from Allocated, it announces $a$ in Hazards and then reads the element, x from Items[a]. If $x\neq\bot$, it performs t&s() on the associated test&set object, TS[a], and, if successful, it clears its announcement, writes to TakeDone, and returns $x$. If the process completes the for loop, it clears its announcement, and rereads InsertDone. If the value of InsertDone has not changed since its previous read, it writes to TakeDone and returns EMPTY. Otherwise, it repeats the entire sequence of steps. Pseudocode for our implementation appears in Figure 5.
7.1 Strong-Linearizability
Consider any execution consisting of operations on this data structure. We linearize the operations as follows:
  • An Insert operation that obtains false when it reads TakeDone on LABEL:b-bounded:insert:reread_TakeDone is linearized at this read. It returns FULL.
  • Consider an Insert(x) operation, insins, that allocated location $a$ to m on LABEL:b-bounded:insert:choose_index and wrote $x$ to Items[a] on LABEL:b-bounded:insert:write_item. Suppose that, while it is poised to write to InsertDone on LABEL:b-bounded:insert:write_InsertDone, some Take operation, $tk$, performs a successful t&s() on TS[a] and then some (possibly different) Take operation writes to TakeDone. In this case, insins is linearized immediately before $tk$, insins returns OK, and we say that insins and $tk$ are coupled.
    We prove in Lemma 14 below that at most one successful t&s() can be performed on TS[a] after insins writes to Items[a] and before it writes to InsertDone. Hence, an Insert operation is coupled with at most one Take operation.
  • Suppose that an Insert operation, insins, writes to InsertDone on LABEL:b-bounded:insert:write_InsertDone, but has not already been linearized as a coupled operation. Then, insins is linearized when it performs this write and returns OK, and we say that insins is uncoupled.
  • Consider a Take operation, $tk$, that performs a successful t&s() on TS[a] on LABEL:b-bounded:take:t&s. It is linearized at the first among the following events to occur after the successful t&s() by $tk$: a write to TakeDone by any Take operation on LABEL:b-bounded:take:successful_writes_TakeDone or LABEL:b-bounded:take:failing_writes_TakeDone, a read of 1 from TS[a] on LABEL:b-bounded:insert:check_TS, and a write to InsertDone on LABEL:b-bounded:insert:write_InsertDone by an uncoupled Insert that wrote to Items[$a^{\prime}$] for some $a^{\prime}\neq a$. Multiple successful Take operations linearized at the same step are ordered arbitrarily and before any other operations linearized at this step, with one exception: a coupled Insert is linearized right before its coupled Take operation. In all cases, $tk$ returns the value it last read on LABEL:b-bounded:take:read_item.
  • Consider a Take operation, $tk$, that obtains false when it reads InsertDone on LABEL:b-bounded:take:reread_InsertDone. If no uncoupled Insert operation writes to InsertDone while $tk$ is poised to write to TakeDone, then $tk$ is linearized at its write to TakeDone (after any other operations linearized at this step). Otherwise, $tk$ is linearized at the first such write to InsertDone, before the Insert and after any successful Take operations linearized at this step. The ordering among unsuccessful Take operations linearized at this write is arbitrary. In both cases, $tk$ returns EMPTY.
To make the linearization points clearer, we also list the types of steps at which operations are linearized. For each, we specify the operations that can be linearized there. If multiple operations can be linearized at the same step, we specify the ordering of these operations.
  • An Insert reads 1 from TS[m] on LABEL:b-bounded:insert:check_TS:
    • A successful Take that previously performed a successful t&s on TS[m].
  • A successful uncoupled Insert, insins, which wrote to Items[m], writes to InsertDone on LABEL:b-bounded:insert:write_InsertDone:
    • Successful Takes that previously performed a successful t&s on TS[m’] for $m^{\prime}\neq m$, arbitrarily ordered.
    • Unsuccessful Takes, arbitrarily ordered.
    • insins.
  • An unsuccessful Insert, insins, obtains false from TakeDone.dRead() on LABEL:b-bounded:insert:reread_TakeDone:
    • insins.
  • A successful Take writes to TakeDone on LABEL:b-bounded:take:successful_writes_TakeDone:
    • Successful Takes, arbitrarily ordered.
        Successful coupled Inserts, each immediately before its coupled Take.
  • An unsuccessful Take $tk$ writes to TakeDone on LABEL:b-bounded:take:failing_writes_TakeDone:
    • Successful Takes, arbitrarily ordered.
        Successful coupled Inserts, each immediately before its coupled Take.
    • tktk.
Next, we provide some intuition explaining our choice of linearization points, as well as why unsuccessful Take operations write to TakeDone on LABEL:b-bounded:take:failing_writes_TakeDone before returning EMPTY.
We pick the linearization point for an Insert operation as we do in Section 6: a successful uncoupled Insert is linearized at its write to InsertDone, a successful coupled Insert operation is linearized right before its coupled Take operation, and an unsuccessful Insert operation is linearized when it obtains false from TakeDone.dRead() on LABEL:b-bounded:insert:reread_TakeDone.
We want to linearize a successful Take operation at its write to TakeDone, which is intended to notify the producer that the bag is not full, similarly to how a successful uncoupled Insert operation notifies the consumers that the bag is not empty. However, in the following four situations, we linearize a successful Take operation earlier than its write to TakeDone.
Suppose a successful Take operation, $tk$, has performed a successful t&s operation on LABEL:b-bounded:take:t&s, but has not yet been linearized when another successful Take operation writes to TakeDone. Then $tk$ is linearized at this write. Linearizing $tk$ at the earliest possible write to TakeDone in this case is done to simplify the linearization.
An unsuccessful Take operation, $tk$, cannot be linearized when it obtains false from InsertDone.dRead(), because there might be successful Take operations that remove elements from the bag, but are not yet linearized. These operations must be linearized before $tk$. To ensure this, $tk$ writes to TakeDone before returning EMPTY. This linearizes every successful Take operation that has performed a successful t&s operation on LABEL:b-bounded:take:t&s, but has not been linearized before the write to TakeDone by $tk$. In this case, $tk$ is linearized at this write, after all such successful Take operations.
A coupled Take operation witnesses that an Insert operation has happened and causes it to be linearized. Similarly, an Insert operation, insins, that witnesses a successful Take operation causes it to be linearized: If insins reads 1 from TS[a] on LABEL:b-bounded:insert:check_TS, but the Take that last performed a successful TS[a].t&s() is not yet linearized, then the Take is linearized at this read, which is before insins removes $a$ from alloc on LABEL:b-bounded:insert:remove_from_alloc. Then, when insins sees that |alloc| < $b$ on LABEL:b-bounded:insert:alloc<b, the bag is not full and it may insert a new element.
Suppose an unsuccessful Take operation, $tk$, got false from InsertDone.dRead(), but has not yet written to TakeDone when an uncoupled Insert operation, insins, writes to InsertDone on LABEL:b-bounded:insert:write_InsertDone. Then $tk$ must be linearized while the bag is still empty. Hence $tk$ is linearized before insins, at this dWrite. Each successful Take operation that performed a successful t&s on LABEL:b-bounded:take:t&s, but has not yet been linearized, is also linearized at this step, before $tk$, to ensure the bag is empty when $tk$ is linearized. There is one exception: If insins wrote to Items[a] on LABEL:b-bounded:insert:write_item, a Take operation that performed a successful TS[a].t&s() at the same location after this write should be linearized after insins. A Take operation that performed a successful TS[a].t&s() at the same location before this write is guaranteed to be linearized before insins. This is because it is linearized at or before the producer last read 1 from TS[a] on LABEL:b-bounded:insert:check_TS, which occurs before the producer removes location $a$ from alloc on LABEL:b-bounded:insert:remove_from_alloc, which, in turn, must occur before insins is allocated location $a$ on LABEL:b-bounded:insert:choose_index.
We proceed to prove that the algorithm in Figure 5 is a strongly-linearizable implementation of a $b$-bounded bag with a single producer.
Fix an execution consisting of possibly concurrent operations in Figure 5. Each operation $op$ in the execution is linearized either when a step by $op$ occurs, or when a step by another operation occurs between two steps of $op$. Thus, the linearization point defined above for every operation is within its interval. Each operation is linearized at the same point in every extension of the execution, hence, if the implementation is linearizable, it is strongly-linearizable.
We now prove a number of useful facts about the implementation.
Items is only modified by the producer. It uses its local set alloc to keep track of the locations in Items that store elements.
Lemma 10
If Items[$a$] $\neq\bot$, then $a\in$ alloc.
Proof 10
Before the producer writes an element $x$ to Items[m] on LABEL:b-bounded:insert:write_item, it adds m to alloc on LABEL:b-bounded:insert:add_to_alloc. Immediately after the producer writes $\bot$ to Items[m] on LABEL:b-bounded:insert:clear_item, it removes m from alloc on LABEL:b-bounded:insert:remove_from_alloc.
When a location in TS is reset on LABEL:b-bounded:insert:reset_TS, this location is in used, which is a set that is local to the producer. We show that this implies that the corresponding location in Items contains $\bot$.
Lemma 11
If $a\in$ used, then Items[$a$] = $\bot$.
Proof 11
The producer adds the location m to used on LABEL:b-bounded:insert:add_to_used, after it sets Items[m] to $\bot$ on LABEL:b-bounded:insert:clear_item. When the producer writes an element $x$ to Items[m] on LABEL:b-bounded:insert:write_item, m is not in used. This is because used is a subset of hazardous, from LABEL:b-bounded:insert:narrow_used_down, and m s not in hazardous, from LABEL:b-bounded:insert:choose_index.
When a Take operation successfully returns an element, we can identify an interval immediately before it performed its successful t&s() in which this test&set object was 0 and the corresponding location in Items contained the returned value.
Lemma 12
Suppose $tk$ is a Take operation that performs a successful t&s() of TS[$a$] on LABEL:b-bounded:take:t&s. Let $x$ be the element that $tk$ last read from Items[$a$] on LABEL:b-bounded:take:read_item. Then TS[$a$] = 0 and Items[$a$] = $x$ between when the producer last wrote to Items[$a$] prior to the read and when $tk$ performed its successful t&s().
Proof 12
Let $i$ be the index of the process that performs $tk$. Prior to reading $x$ from Items[$a$], tktk wrote $a$ to Hazards[$i$] on LABEL:b-bounded:take:add_hazard. It next writes to Hazards[$i$] on LABEL:b-bounded:take:clear_hazard1. So Hazards[$i$] = $a$ between the step $s_{1}$ when $tk$ last performed LABEL:b-bounded:take:read_item and the step $s_{2}$ when it last performed LABEL:b-bounded:take:t&s.
Note that TS[$a$] = 0 immediately before $tk$ performed its successful t&s() on TS[$a$]. Let $s_{0}$ be the step when the producer last wrote to Items[$a$] (on LABEL:b-bounded:insert:write_item) prior to $s_{1}$. So, Items[$a$] = $x$ between $s_{0}$ and $s_{1}$. Assume, for the sake of contradiction, that TS[$a$] = 1 at some configuration between $s_{0}$ and $s_{2}$. Then between this configuration and $s_{2}$, the producer must have reset TS[$a$] on LABEL:b-bounded:insert:reset_TS. From LABEL:b-bounded:insert:reset_loop, $a$ was in used when the reset was performed, so, by Lemma 11, Items[$a$] = $\bot$ when the reset was performed. Thus, the producer must have written $\bot$ to Items[$a$] on LABEL:b-bounded:insert:clear_item before it performed the reset. Since Items[$a$] =x=x between $s_{0}$ and $s_{1}$, the producer must have written $\bot$ to Items[$a$], collected Hazards on LABEL:b-bounded:insert:collect_hazards, and performed the reset between $s_{1}$ and $s_{2}$. In particular, the producer reads $a$ from Hazards[i]. But, by LABEL:b-bounded:insert:reset_loop, the producer does not reset TS[$a$] on LABEL:b-bounded:insert:reset_TS. This is a contradiction. Thus TS[$a$] = 0 between $s_{0}$ and $s_{2}$.
By the test on LABEL:b-bounded:insert:check_TS, the producer does not write $\bot$ to Items[$a$] on LABEL:b-bounded:insert:clear_item and $a$ is not removed from alloc on LABEL:b-bounded:insert:remove_from_alloc between $s_{0}$ and $s_{2}$. Thus, the producer does not choose the index $a$ again on LABEL:b-bounded:insert:choose_index and does not write to Items[$a$] on LABEL:b-bounded:insert:write_item between these two steps. Hence, Items[$a$] = $x$ between when the producer last wrote to Items[$a$] at step $s_{0}$ and when $tk$ performed its successful t&s() at step $s_{2}$.
Next, we show that when the producer writes an element into some location in Items, that element is available to be taken by a taker.
Lemma 13
When the producer writes an element to Items[m] on LABEL:b-bounded:insert:write_item, TS[m] = 0.
Proof 13
Consider a step $s$ of the execution in which the producer performs a write to TS[$a$] on LABEL:b-bounded:insert:write_item. All locations of Items initially contain $\bot$ and elements are only written into Items on LABEL:b-bounded:insert:write_item. Thus, if $s$ is the first step in which the producer performs a write to TS[$a$] on LABEL:b-bounded:insert:write_item, then Items[a] = $\bot$ prior to $s$.
Suppose that, prior to $e$, the producer has not performed LABEL:b-bounded:insert:write_item with m = $a$. All locations of Items initially contain $\bot$, non-$\bot$ values are only written into Items on LABEL:b-bounded:insert:write_item, and t&s() on TS[a] is only performed on LABEL:b-bounded:take:t&s after a taker sees that Items[a] $\neq\bot$, no t&s() on TS[a] was performed prior to $s$. Since TS[a] is initially 0, it is 0 when $s$ is performed.
Otherwise, let $s^{\prime}$ be the last step prior to $s$ at which the producer performed LABEL:b-bounded:insert:write_item with m = $a$. Then $a$ was added to alloc when the producer last performed LABEL:b-bounded:insert:add_to_alloc prior to $s^{\prime}$. When the producer last performs LABEL:b-bounded:insert:choose_index prior to $s$, the location $a$ was not in alloc. Thus, between these two steps, the producer removed $a$ from alloc on LABEL:b-bounded:insert:remove_from_alloc and added $a$ to used on LABEL:b-bounded:insert:add_to_used.
By Lemma 11, Items[a] = $\bot$ while $a\in$ used. Immediately after step $s$, Items[$a$] $\neq\bot$, so $a\not\in$ used. Prior to being removed from used on LABEL:b-bounded:insert:narrow_used_down, $a$ was not in hazardous, so TS[a] was reset to 0 on LABEL:b-bounded:insert:reset_TS. Since $a\in$ used when this step occurred, Items[$a$] = $\bot$. Between this step and $s$, there is no write to Items[$a$] on LABEL:b-bounded:insert:write_item, by definition of $s^{\prime}$, so Items[$a$] = $\bot$. When a successful t&s() on TS[$a$] is performed, Lemma 12 implies that Items[a] $\neq$ $\bot$. Thus, no successful t&s() on TS[$a$] was performed between these two steps. Hence, TS[$a$] = 0 when $s$ was performed.
The next lemma says that an Insert operation is coupled with at most one Take operation.
Lemma 14
Consider an Insert(x) operation, insins, that allocated location $a$ to m on LABEL:b-bounded:insert:choose_index. At most one successful t&s() on TS[$a$] by a Take operation is performed while insins is poised to write to InsertDone on LABEL:b-bounded:insert:write_InsertDone.
Proof 14
By Lemma 13, when insins writes to Items[a] on LABEL:b-bounded:insert:write_item, TS[a] = 0. Assume some Take operation performs a successful t&s() on TS[a] after insins performs this write and before it performs its next step, which is a write to InsertDone on LABEL:b-bounded:insert:write_InsertDone. For another Take operation to perform a successful t&s() on TS[a], TS[a] must be first reset to 0. It is reset only by the producer (on LABEL:b-bounded:insert:reset_TS). Therefore, no other Take operation can perform a successful t&s() on TS[a] while insins is poised to write to InsertDone.
Between a successful t&s() and a reset of TS[a], the producer reads 1 from TS[a].
Lemma 15
After a Take operation performs a successful TS[a].t&s() the producer must read 1 from TS[a] before it resets TS[a].
Proof 15
Let $ts$ be a successful TS[$a$].t&s() performed by a Take operation on LABEL:b-bounded:take:t&s. By Lemma 12, when $ts$ is performed, Items[$a$] $\neq\bot$ and, hence, by Lemma 11, $a\notin$ used. When a reset of TS[$a$] is performed on LABEL:b-bounded:insert:reset_TS, the condition on LABEL:b-bounded:insert:reset_loop ensures that $a\in$ used. So $a$ is added to used on LABEL:b-bounded:insert:add_to_used between $ts$ and the next reset of TS[$a$]. Before $a$ is added to used, the producer reads 1 from TS[a] on LABEL:b-bounded:insert:check_TS.
Only the producer performs resets, but does not perform any resets between performing LABEL:b-bounded:insert:check_TS and next performing LABEL:b-bounded:insert:add_to_used. Since TS[a] =0=0 when $ts$ is performed, $ts$ does not occur between these two steps by the producer. Hence, the producer reads 1 from TS[a] after $ts$ and before it next resets TS[$a$].
Two successful Take operations that return elements from the same location in Items are linearized at different steps.
Lemma 16
Two successful Take operations that both perform a successful t&s() on TS[$a$] are not linearized at the same step.
Proof 16
Suppose that $tk_{1}$ is a Take operation that performs a successful t&s(), $ts_{1}$, on TS[$a$] before the Take operation $tk_{2}$ performs a successful t&s(), $ts_{2}$, on TS[$a$]. A reset of TS[$a$] occurs between $ts_{1}$ and $ts_{2}$. By Lemma 15, the producer reads 1 from TS[a] (on LABEL:b-bounded:insert:check_TS) between $ts_{1}$ and performing this reset. Note that $tk_{1}$ is linearized at or before this read. However, $tk_{2}$ is linearized after $ts_{2}$. Thus, they are not linearized at the same step.
The following lemma says that, after an element is written to Items[$a$], the next Take operation to perform a successful t&s() on TS[$a$] returns this element.
Lemma 17
Consider an Insert($x_{1}$) operation that writes $x_{1}$ to Items[$a$] on LABEL:b-bounded:insert:write_item. The Take operation that next performs a successful TS[$a$].t&s() returns $x_{1}$.
Proof 17
Let $tk$ be the Take operation that performs the first successful TS[a].t&s(), $ts$, following the write, $w_{1}$, of $x_{1}$ to Items[$a$]. Let $r$ be the last read of Items[a] on LABEL:b-bounded:take:read_item performed by $tk$ prior to $ts$ and let $x$ be the element that it read during $r$. Let $w$ be the last write to Items[a] on LABEL:b-bounded:insert:write_item by the producer before $r$, which implies this is a write of $x$. Between $w$ and the producer’s next write to Items[a], the producer executes LABEL:b-bounded:insert:choose_index, and at that point $a\notin$ alloc, which implies by Lemma 10 that Items[a] = $\bot$. But by Lemma 12, Items[a] = $x$ between $w$ and $ts$. Therefore, the producer’s next write to Items[a] after $w$ must happen after $ts$. In other words, $w$ was the last write to Items[a] before $ts$, namely, $w=w_{1}$. Hence, $x=x_{1}$. So $tk$ would return $x_{1}$.
To show that the ordering of operations that we defined is a linearization, we prove that the values returned by the operations are consistent with the sequential specifications of a $b$-bounded bag. We do this by induction on the length of the execution.
Let $S$ be the ordering of a prefix, $\sigma$, of the execution, let $e$ be the next step of the execution, and $S^{\prime}$ be the ordering of $\sigma\cdot e$. Suppose $S$ is a linearization. Let $B$ be the set of items in the bag at the end of $S$. Further suppose that for every Take operation $tk$ that appears in $S$, returns an item $x$, and performs in $\sigma$ a successful t&s() on TS[a] for some index $a$, an Insert($x$) operation that wrote in $\sigma$ to Items[a] appears in $S$ before $tk$ and no Take operation returning $x$ appear in $S$ between this Insert operation and $tk$. Due to our no repetitions assumption, this informally means that every Take operation that performs in $\sigma$ a successful t&s() on some location in TS, takes the item that was inserted to the bag by an Insert operation that wrote to the same location in Items. Let $B$ be the set of items in the bag at the end of $S$. We will prove that $S^{\prime}$ is a linearization, and that for every Take operation $tk$ that appears in $S^{\prime}$ but not in $S$, returns an item $x$ and performs in $\sigma$ a successful t&s() on TS[a] for some index $a$, an Insert($x$) operation that wrote in $\sigma$ to Items[a] appears in $S$, such that no Take operation returning $x$ appears in $S$ after it. The latter part is proven in Lemma 21.
It suffices to consider the steps at which operations are linearized, namely when a Take operation writes to TakeDone and when an Insert operation reads TakeDone, writes to InsertDone or reads 1 from TS.
In what follows we will use the following lemmas. The first one states that between two successful Insert operations that write to the same location, a single Take operation that takes from that location is linearized and it returns the item inserted by the first Insert among the two.
Lemma 18
Let $ins_{1}$ and $ins_{2}$ be two successful Insert operations that were allocated the same location $a$. Assume $ins_{1}$ was the first among the two to be executed by the producer. Then a Take operation $tk$ that performs a successful TS[a].t&s() and returns the item inserted by $ins_{1}$ is linearized after $ins_{1}$. If $ins_{2}$ is linearized, then it is linearized after $tk$.
Proof 18
Let $ins_{2}^{\prime}$ be the first successful Insert operation to be executed after $ins_{1}$ among those allocated location $a$. Let $w_{1}$ be the write event to Items[a] by $ins_{1}$ on LABEL:b-bounded:insert:write_item, and $w_{2}$ be the write event to Items[a] by $ins_{2}^{\prime}$. Let $w_{InsertDone}$ be the write event to InsertDone by $ins_{1}$ on LABEL:b-bounded:insert:write_InsertDone, which is its next step after writing to Items[a]. From the code, it is evident that $a\in$ alloc when $w_{InsertDone}$ occurs. Before $ins_{2}^{\prime}$ writes to Items[a], it allocates location $a$ to m on LABEL:b-bounded:insert:choose_index, and at that point $a\notin$ alloc. Therefore, between $w_{InsertDone}$ and $w_{2}$, $m$ is removed from alloc. This is done by the producer on LABEL:b-bounded:insert:remove_from_alloc, after reading 1 from TS[a] on LABEL:b-bounded:insert:check_TS. Let $r_{TS}$ be this read event. Then $r_{TS}$ occurs between $w_{InsertDone}$ and $w_{2}$. By Lemma 13, TS[a] = 0 at $w_{1}$. Thus, between $w_{1}$ and $r_{TS}$, a Take operation performed a successful TS[a].t&s(). Let $tk$ be the Take operation to perform the first such t&s(), and let $ts$ be this t&s() event. $tk$ returns the item inserted by $ins_{1}$ due to Lemma 17.
If $ins_{2}^{\prime}$ is linearized, it is linearized after $w_{2}$. $tk$ is linearized at the latest at $r_{TS}$. Therefore, $tk$ is linearized before $ins_{2}^{\prime}$. To show that $tk$ is linearized after $ins_{1}$, we consider two cases. First consider the case in which when $ins_{1}$ is poised to write to InsertDone on LABEL:b-bounded:insert:write_InsertDone, $ts$ occurs and then some Take operation writes to TakeDone. Then $ins_{1}$ is coupled with $tk$ and linearized before it. Otherwise, $ins_{1}$ is linearized at $w_{InsertDone}$, and $tk$ is linearized at one of the following events, all occurring after $w_{InsertDone}$: a read of 1 from TS[a] by the producer on LABEL:b-bounded:insert:check_TS, a write to InsertDone by an uncoupled Insert operation that is allocated an index $\neq a$, or a write to TakeDone.
The following lemma clarifies the linearization order among a successful Take operation and a successful Insert operation that operate on the same location.
Lemma 19
Let $tk$ be a Take operation that performs a successful TS[a].t&s(), in an event $ts$. Let insins be an Insert operation that writes to Items[a] on LABEL:b-bounded:insert:write_item, in an event $w$ that occurs after $ts$. Then $tk$ is linearized before insins.
Proof 19
By Lemma 13, TS[a] = 0 at $w$. Hence, between $ts$ and $w$, a reset of TS[a] on LABEL:b-bounded:insert:reset_TS must take place. By Lemma 15, between $ts$ and this reset, the producer reads 1 from TS[a] on LABEL:b-bounded:insert:check_TS. $tk$ is linearized at this read if it is not already linearized earlier. Therefore, $tk$ is linearized before $w$, and thus, before insins.
The following lemma states that from the successful t&s() a Take operation performs on TS[a] until this operation is linearized, no Take operation performs a successful TS[a].t&s().
Lemma 20
Let $tk$ be a Take operation that appears in $S^{\prime}$ but not in $S$ and performs in $\sigma$ a successful t&s() on TS[a] for some index $a$. Then no successful TS[a].t&s() occurs in $\sigma$ after $tk$’s successful TS[a].t&s().
Proof 20
Let $ts$ be $tk$’s successful TS[a].t&s(). Before another successful TS[a].t&s() occurs in $\sigma$ after $ts$, a reset of TS[a] must be executed. By Lemma 15, between $ts$ and this reset, the producer reads 1 from TS[a] on LABEL:b-bounded:insert:check_TS. But no such read is executed in $\sigma$ after $ts$, because if it were, $tk$ would be linearized at this step at the latest and would appear in $S$. is executed in $\sigma$ after $ts$. We have that a successful TS[a].t&s() that follows $ts$ must occur after a reset, which in turn occurs after read of 1 from TS[a], which does not occur in $\sigma$ after $ts$. So such a successful t&s() does not occur in $\sigma$.
The following lemma proves the second part of the induction statement. After this lemma we prove the first part, namely, that $S^{\prime}$ is a linearization.
Lemma 21
Let $tk$ be a Take operation that appears in $S^{\prime}$ but not in $S$, returns an item $x$ and performs in $\sigma$ a successful t&s() on TS[a] for some index $a$. Let insins be the last Insert operation to write to Items[a] in $\sigma$ before $tk$ read it for the last time (which it did before executing its successful t&s()). If insins appears in $S$, then no Take operation that returns $x$ appears after it in $S$.
Proof 21
Let $ts$ be $tk$’s successful TS[a].t&s(). Let $w$ be the write to Items[a] on LABEL:b-bounded:insert:write_item by insins. We first consider Take operations in $S$ that perform in $\sigma$ a successful t&s() on TS[a], if such operations exist. No successful t&s() is performed on TS[a] between $w$ and $ts$, as TS[a] = 0 in this interval by Lemma 12. No successful t&s() is performed on TS[a] in $\sigma$ after $ts$ by Lemma 20. Any Take operation performing a successful TS[a].t&s() before $w$, appears in $S$ before insins by Lemma 19.
Now, we consider Take operations in $S$ returning $x$, which perform in $\sigma$ a successful t&s() on TS[a’] for some index $a^{\prime}\neq a$. We will show that if any such operations exist, they must appear before insins in $S$. Assume for sake of contradiction that such an operation appears after insins in $S$. Let $tk^{\prime}$ be the first such operation to appear in $S$ after insins. Assume it performed a successful t&s() on TS[a’] for some location $a^{\prime}$. Due to the induction hypothesis, an Insert($x$) operation that wrote in $\sigma$ to Items[a’] appears in $S$ before $tk^{\prime}$, such that $tk^{\prime}$ is the first Take operation returning $x$ to appear after this Insert($x$). This Insert operation is not insins, because they write to different locations in Items. $tk^{\prime}$ is also the first Take operation returning $x$ to appear in $S$ after insins, due to the definition of $tk^{\prime}$ in addition to what we earlier proved—that Take operations that perform in $\sigma$ a successful t&s() on TS[a] do not appear in $S$ after insins. Hence, right before $tk^{\prime}$, $x$ appears at least twice in the bag. This is a contradiction to our no repetitions assumption.
First, suppose $e$ is an execution of TakeDone.dRead() on LABEL:b-bounded:insert:reread_TakeDone by an Insert operation insins, which returns false. Then $S^{\prime}$ is $S$ followed by insins with return value FULL. To show that $S^{\prime}$ is a linearization, we must show that $|B|=b$.
Lemma 22
$|alloc|\leq b$.
Proof 22
alloc is initialized to $\phi$. A location is added to alloc only on LABEL:b-bounded:insert:add_to_alloc, when $|alloc|<b$.
Let $C$ be the configuration at the end of $\sigma$, let $C^{*}$ be the configuration immediately after the last previous read of TakeDone by insins on either Line 33 or LABEL:b-bounded:insert:reread_TakeDone, let $S^{*}$ be the linearization of the prefix of $\sigma$ ending with $C^{*}$, and let $B^{*}$ be the set of items in the bag at the end of $S^{*}$.
$|alloc|=b$ when insins executes LABEL:b-bounded:insert:alloc<b after $C^{*}$: $|alloc|$ cannot be smaller than $b$ at that step because insins continues to the else if clause, and it cannot be bigger than $b$ due to Lemma 22. insins does not execute LABEL:b-bounded:insert:remove_from_alloc between $C^{*}$ and $C$, because remove a location from alloc would imply that $|alloc|>b$ before the removal, which is impossible by Lemma 22. So insins does not modify allocalloc between $C^{*}$ and $C$. Let $alloc^{*}$ be the value of allocalloc in this interval. We have $|alloc^{*}|=b$. This implies that insins obtains 0 from the read of TS[m] for all indices $m\in alloc$ in its executions of LABEL:b-bounded:insert:check_TS following $C^{*}$.
As insins gets false from TakeDone.dRead() at $e$, no write to TakeDone is performed between $C^{*}$ and $C$. This implies that no successful Take operations are linearized between these configuration, as these are linearized either at a write to TakeDone or at steps by the producer which are not carried out in this interval (reading 1 from TS[m] on LABEL:b-bounded:insert:check_TS or writing to InsertDone on LABEL:b-bounded:insert:write_InsertDone). So to prove that $|B|=b$, it is enough to show that $|B^{*}|=b$.
Let $m$ be a location in $alloc^{*}$. We will show that the item that is in Items[m] at $C^{*}$ is in $B^{*}$. Since $|alloc^{*}|=b$ and due to our no repetitions assumption, this will imply $|B^{*}|=b$. Let $ins^{\prime}$ be the last Insert operation to add $m$ to alloc on LABEL:b-bounded:insert:add_to_alloc before $C^{*}$. $ins^{\prime}$ is successful so $ins^{\prime}\neq ins$. Hence, $ins^{\prime}$ was completed before $C^{*}$, and appears in $S^{*}$ as a successful insertion. We will show that a Take operation that takes the item inserted by $ins^{\prime}$ cannot appear in $S^{*}$ after $ins^{\prime}$.
Let $w$ be the write to Items[m] on LABEL:b-bounded:insert:write_item by $ins^{\prime}$. At $w$, $m\notin$ hazardous according to LABEL:b-bounded:insert:choose_index, while used $\subseteq$ hazardous by LABEL:b-bounded:insert:narrow_used_down. Therefore, $m$ $\notin$ used at $w$.After $w$, $m$ is added to used on LABEL:b-bounded:insert:add_to_used only after it is removed from alloc on LABEL:b-bounded:insert:remove_from_alloc. The latter may happen only after $C$: it cannot happen until $C^{*}$ by definition of $ins^{\prime}$, and it cannot happen between $C^{*}$ until $C$ since alloc remains unchanged in this interval. Thus, $m$ $\notin$ used continues to hold from $w$ until $C$. Therefore, no reset of TS[m] is executed on LABEL:b-bounded:insert:reset_TS between $w$ and $C$. Recall that $ins^{\prime}$ reads TS[m] between $C^{*}$ and $C$ and obtains 0. As TS[m] is not reset between $w$ and this read, a successful TS[m].t&s() cannot be executed within this interval, specifically not between $w$ and $C^{*}$. A Take operation that executes a successful TS[m].t&s() before $w$ is not linearized after $ins^{\prime}$ by Lemma 19, and a Take operation that executes a successful TS[m].t&s() after $C^{*}$ is not linearized at $S^{*}$. Therefore, no Take operation is linearized after $ins^{\prime}$ in $S^{*}$.
Next, suppose $e$ is an execution of a read of 1 from TS[a] by an Insert operation insins on LABEL:b-bounded:insert:check_TS. Further suppose a Take operation, $tk$, performed a successful t&s() on TS[a] on LABEL:b-bounded:take:t&s, in a step $ts$. Assume between $ts$ and $e$, none of the following occurred: another read of 1 from TS[a] by the producer, a write to TakeDone, or a write to Items[a’] for $a^{\prime}\neq a$ by an uncoupled Insert operation. There could not be more than one such Take operation by Lemma 16. Let the item $tk$ obtained from Items[a] in its last execution of LABEL:b-bounded:take:read_item be $x$. Then $S^{\prime}$ is $S$ followed by $tk$ with return value $x$. To show that $S^{\prime}$ is a linearization, we must show that $x\in B$.
Let $ins^{\prime}$ be the last Insert operation to write to Items[a] on LABEL:b-bounded:insert:write_item before $tk$ reads it for the last time (which it does before $ts$), and let $w$ be this write. Then $w$ is a write of $x$ and $ins^{\prime}$ in an Insert($x$) operation. $ins^{\prime}\neq ins$ since $ins^{\prime}$ does not execute LABEL:b-bounded:insert:check_TS after $w$. Therefore, $ins^{\prime}$ is completed in $\sigma$ and thus appears in $S$. Since $S$ is a linearization, in the prefix of $S$ that ends with $ins^{\prime}$, the bag contains $x$. By Lemma 21, there is no Take operation that returns $x$ in $S$ after $ins^{\prime}$. This implies $x\in B$.
Now, suppose $e$ is an execution of InsertDone.dWrite() by an uncoupled Insert($x$) operation insins on LABEL:b-bounded:insert:write_InsertDone. Let $m$ be the index allocated to insins when it executed LABEL:b-bounded:insert:choose_index. Let $tk_{i}$ for $1\leq i\leq s$ denote every Take operation that performed in $\sigma$ a successful t&s() on TS[$a_{i}$] where $a_{i}\neq m$ such that since the t&s() until the end of $\sigma$, no write to TakeDone was executed, no write to InsertDone by an uncoupled Insert operation was executed and no read of 1 from TS[$a_{i}$] by an Insert operation was executed. Let the item $tk_{i}$ obtained from Items[$a_{i}$] in its last execution of LABEL:b-bounded:take:read_item be $x_{i}$. Let $f$ be the number of Take operations that obtained false from InsertDone.dRead() on LABEL:b-bounded:take:reread_InsertDone prior to $e$, and between the read and $e$ there was no write to InsertDone and they have not yet executed TakeDone.dWrite(). $S^{\prime}$ is $S$ followed by $tk_{1},\ldots,tk_{s}$, with return value $x_{1},\ldots,x_{s}$ respectively, in an arbitrary order, then the above-mentioned $f$ failing Take operations in an arbitrary order, all with return value EMPTY, and then insins with return value OK. To show that $S^{\prime}$ is a linearization, we must show that (i) $\{x_{1},\ldots,x_{s}\}\subseteq B$, (ii) $B=\{x_{1},\ldots,x_{s}\}$ if $f>0$, and (iii) $|B|<b$ if $s=f=0$. Due to the no repetitions assumption, (ii) would imply $|B|=s$.
For each $1\leq i\leq s$, let $ins_{i}$ be the Insert operation that performed the last write to Items[$a_{i}$] before $tk_{i}$ read it for the last time (which it did before executing its successful t&s()). Hence, $ins_{i}$ in an Insert($x_{i}$) operation. We begin with proving (i). Let $i$ be an index such that $1\leq i\leq s$. $ins\neq ins_{i}$ since insins was allocated the index $m$ while $ins_{i}$ was allocated the index $a_{i}\neq m$. Since $e$ is a step performed by insins and there is a single producer, $ins_{i}$ was completed in $\sigma$. Hence $ins_{i}$ appears in $S$. Since $S$ is a linearization, in the prefix of $S$ that ends with $ins_{i}$, the bag contains $x_{i}$. By Lemma 21, there is no Take operation after $ins_{i}$ in $S$ that takes $x_{i}$. This implies $x_{i}\in B$.
We proceed to prove (ii). Let $tk$ be one of the $f$ Take operations that obtained false from InsertDone.dRead() on LABEL:b-bounded:take:reread_InsertDone prior to $e$ and have not yet executed TakeDone.dWrite(). We will prove that no items other than $x_{i}$ for $1\leq i\leq s$ are in $B$.
Let $C$ be the configuration at the end of $\sigma$ (immediately before $e$), let $\widehat{C}$ be the configuration immediately after $tk$ obtained false from InsertDone.dRead() on LABEL:b-bounded:take:reread_InsertDone, let $C^{*}$ be the configuration immediately after the last previous read of InsertDone by $tk$ on either LABEL:b-bounded:take:read_InsertDone or LABEL:b-bounded:take:reread_InsertDone, let $S^{*}$ be the linearization of the prefix of $\sigma$ ending with $C^{*}$, and let $B^{*}$ be the set of items in the bag at the end of $S^{*}$.
Considering $tk$ obtained false from InsertDone.dRead() right before $\widehat{C}$, the producer does not write to InsertDone between $C^{*}$ and $\widehat{C}$. No write to InsertDone is performed by an uncoupled Insert operation between $\widehat{C}$ and $C$ as $e$ is the first such write since $\widehat{C}$.
Suppose there was a successful Insert operation that is in $S$, but not in $S^{*}$, which means it is linearized between configurations $C^{*}$ and $C$. Since no uncoupled Insert operation wrote to InsertDone on LABEL:b-bounded:insert:write_InsertDone between $C^{*}$ and $C$, this Insert operation was coupled with a Take operation. Hence it was immediately taken from the bag after it was inserted into the bag. Therefore, any item in $B$ was inserted to the bag in $S^{*}$ and was in $B^{*}$.
So, suppose there is $x^{*}\in B^{*}$ such that $x^{*}\neq x_{i}$ for all $1\leq i\leq s$. Items inserted by coupled Insert operations are immediately taken from the bag by a Take operation linearized at the same step, so they are not in the bag in the end of the linearization of any prefix of the execution. Hence they are not in $B^{*}$. Therefore, the item $x^{*}$ was inserted into the bag by an uncoupled Insert operation, $ins^{*}$. (Each item in the bag may be uniquely associated with the Insert operation that inserted it thanks to our no repetitions assumption.) This operation wrote $x^{*}$ to Items in some location, $a^{*}$, on LABEL:b-bounded:insert:write_item, and then wrote to InsertDone on LABEL:b-bounded:insert:write_InsertDone before $C^{*}$. We will consider three distinct cases: (1) $a^{*}=a_{i}$ for some $1\leq i\leq s$, (2) $a^{*}=m$, and (3) $a^{*}\neq a_{i}$ for all $1\leq i\leq s$ and $a^{*}\neq m$. For each case, we will show that a Take operation returning $x^{*}$ appears in $S$ after $ins^{*}$.
We start with case (1). $tk_{i}$ read $x_{i}\neq x^{*}$ from Items[$a^{*}$] on LABEL:b-bounded:take:read_item. We showed in the proof of (i) that $ins_{i}$ appears in $S$. If $ins_{i}$ appears in $S$ before $ins^{*}$, then by Lemma 18, a Take operation that performed a successful TS[$a^{*}$].t&s() appears in $S$ after $ins_{i}$. But this is impossible: in the end of the above proof of (i), we showed that no such Take operations appear in $S$ after $ins_{i}$. Therefore, $ins^{*}$ must appear before $ins_{i}$ in $S$. By Lemma 18, a Take operation that takes $x^{*}$ appears in $S$ after $ins^{*}$, and we are done with case (1).
For case (2), $ins^{*}\neq ins$ since the first appears in $S$ and the latter does not. By Lemma 18, a Take operation that takes $x^{*}$ appears in $S$ after $ins^{*}$.
We proceed to case (3). When $ins^{*}$ wrote to Items[$a^{*}$], TS[$a^{*}$] = 0 by Lemma 13. We consider several sub-cases.
(3a) If $a^{*}$ was removed from Allocated on LABEL:b-bounded:insert:add_to_Allocated before $tk$ read Allocated on LABEL:b-bounded:take:read_Allocated after $C^{*}$, then before this read the producer has removed $a^{*}$ from allocalloc on LABEL:b-bounded:insert:remove_from_alloc, which means the producer passed LABEL:b-bounded:insert:check_TS after $ins^{*}$ was completed. (3b) If $tk$ read $a^{*}$ from Allocated on LABEL:b-bounded:take:read_Allocated after $C^{*}$ and then read $\bot$ from Items[$a^{*}$] on LABEL:b-bounded:take:read_item, then the producer wrote $\bot$ to Items[$a^{*}$] on LABEL:b-bounded:insert:clear_item after $ins^{*}$ was completed, and before this write the producer passed LABEL:b-bounded:insert:check_TS. So in both cases—(3a) and (3b), the producer read 1 from TS[$a^{*}$] on LABEL:b-bounded:insert:check_TS after $ins^{*}$ was completed. Since TS[$a^{*}$] = 0 when $ins^{*}$ wrote to Items[$a^{*}$], some Take operation performed a successful t&s() on TS[$a^{*}$] after $ins^{*}$ wrote to Items[$a^{*}$] and before $\widehat{C}$. Let $tk^{*}$ be the first Take to perform such a t&s(). It appears in $S$ because it is linearized at the latest at the read of 1 from TS[$a^{*}$] by the producer.
(3c) Otherwise, after $C^{*}$, $tk$ read $a^{*}$ from Allocated and then read a non-$\bot$ value from Items[$a^{*}$]. In this case, $tk$ must have performed an unsuccessful t&s() of TS[$a^{*}$] on LABEL:b-bounded:take:t&s. As TS[$a^{*}$] = 0 when $ins^{*}$ wrote to Items[$a^{*}$], another Take operation performed a successful TS[$a^{*}$].t&s() after $ins^{*}$ wrote to Items[$a^{*}$] and before the unsuccessful t&s() by $tk$. Let $tk^{*}$ be the first Take to perform such a t&s(). It appears in $S$ because otherwise it would be one of $tk_{i}$ for some $1\leq i\leq s$, but it is not as it operates on location $a^{*}\neq a_{i}$ for all $1\leq i\leq s$.
In all cases—(3a), (3b) and (3c), $tk^{*}$ takes $x^{*}$ from the bag by Lemma 17. Additionally, $tk^{*}$ appears in $S$ after $ins^{*}$: $tk^{*}$ is linearized after $ins^{*}$ wrote to Items[$a^{*}$] because $tk^{*}$ performs its successful TS[$a^{*}$].t&s() after that write. Since $ins^{*}$ is uncoupled, $tk^{*}$ cannot be linearized before the write to InsertDone by $ins^{*}$, which is the linearization point of $ins^{*}$.
We proceed to prove (iii). Assume $s=f=0$. We use the following lemma, which states that the items that are in the bag in the end of the linearization of any prefix of $\sigma$ are a subset of the non-$\bot$ items in Items at the end of the prefix.
Lemma 23
For any prefix $\tau$ of $\sigma$, the items in the bag in the end of the linearization of $\tau$ is a subset of the non-$\bot$ items that are in Items at the end of $\tau$.
Proof 23
We will show that when an item is added to the bag, it appears in some location in Items; and that before it is removed from this location in Items it is removed from the bag.
An item inserted to the bag by a coupled Insert operation does not appear in the bag in the end of the linearization of any prefix of $\sigma$, since it is inserted and then taken at the same step. Thus, an item $x$ becomes a part of the bag of the linearization in a step $w_{InsertDone}$ in which a successful uncoupled Insert($x$) operation $ins^{\prime}$ writes to InsertDone and is linearized. Let $a$ be the location that $ins^{\prime}$ is allocated. Let $w$ be the write of $x$ to Items[a] by $ins^{\prime}$ on LABEL:b-bounded:insert:write_item, which happens before $w_{InsertDone}$. Items[a] still contains $x$ in $w_{InsertDone}$, because after $w$, the producer writes to Items[a] only in following Insert operations, after $ins^{\prime}$ is completed.
We proceed to the second part, showing that if $x$ is removed from Items, it is removed from the bag beforehand. If Items[a] = $x$ since $w_{InsertDone}$ until the end of $\sigma$, we are done. Else, let $w_{\bot}$ be the step after which Items[a] $\neq x$ for the first time since $w_{InsertDone}$. When the producer writes $x$ to Items[a] at $w$, $a$ is in alloc after it was added to alloc on LABEL:b-bounded:insert:add_to_alloc. To perform another write to Items[a] on LABEL:b-bounded:insert:write_item, the producer needs to choose the index $a$ again on LABEL:b-bounded:insert:choose_index, which requires that $a$ $\notin$ alloc. Before that, $a$ must be removed from alloc on LABEL:b-bounded:insert:remove_from_alloc. Hence, after writing $x$ to Items[a] at $w$, the producer next writes to Items[a] on LABEL:b-bounded:insert:clear_item. Namely, $w_{\bot}$ is a write of $\bot$ to Items[a] by the producer. We will show that a Take operation takes $x$ from the bag of the linearization, after $ins^{\prime}$ inserts it to the bag of the linearization, in a step that precedes $w_{\bot}$. Let $r_{TS}$ be the producer’s step preceding $w_{\bot}$, in which it executes LABEL:b-bounded:insert:check_TS. TS[a] = 1 at $r_{TS}$. By Lemma 13, TS[a] = 0 at $w$. So a successful TS[a].t&s() is performed by a Take operation between $w$ and $r_{TS}$. Let $tk$ be the Take operation that performs the first such t&s, in a step $ts$. tktk is linearized at the latest at $r_{TS}$, hence it is linearized before $w_{\bot}$. It takes $x$ from the bag by Lemma 17. It remains to show that $tk$ is linearized after $ins^{\prime}$. $tk$ is linearized after $w$ because $tk$ is linearized after $ts$ which happens after $w$. Since $ins^{\prime}$ is uncoupled, $tk$ cannot be linearized between $w$ and $w_{InsertDone}$. Therefore, $tk$ is linearized after $w_{InsertDone}$, which is the linearization point of $ins^{\prime}$.
Let $C$ be the configuration that appears in $\sigma$ right before the addition of $m$ to alloc by insins on LABEL:b-bounded:insert:add_to_alloc. Let $B_{C}$ be the set of items in the bag at the end of the linearization of $\sigma$’s prefix that ends with $C$. By Lemma 23, $|B_{C}|$ is less than or equal to the number of non-$\bot$ items that are in Items at $C$. By Lemma 10, the locations in Items with non-$\bot$ values at $C$ are a subset of the locations in alloc at the same configuration, and there are less than $b$ locations in alloc according to the check on LABEL:b-bounded:insert:alloc<b. Hence, $|B_{C}|<b$. No Insert operations are linearized in $\sigma$ after $C$, since the producer is executing the uncoupled Insert insins in this interval, which is linearized only at $e$ that comes right after $\sigma$. Therefore, for any prefix of $\sigma$ that extends the prefix ending with $C$, the bag in the end of the linearization of this prefix does not have more items than $B_{C}$, and we are done.
Finally, suppose $e$ is a write to TakeDone by a Take() operation $tk$ on LABEL:b-bounded:take:successful_writes_TakeDone or LABEL:b-bounded:take:failing_writes_TakeDone. Let $tk_{i}$ for $1\leq i\leq s$ denote every Take operation that performed in $\sigma$ a successful t&s() on TS[$a_{i}$] such that no Insert operation has written to Items[$a_{i}$] before this t&s() but has not yet written to InsertDone at $e$. Let the item $tk_{i}$ obtained from Items[$a_{i}$] in its last execution of LABEL:b-bounded:take:read_item be $x_{i}$. Let $tk^{\prime}$ be a Take operation that performed in $\sigma$ a successful t&s() on TS[$a^{\prime}$] such that an Insert($x^{\prime}$) operation, $ins^{\prime}$, has written to Items[$a^{\prime}$] before this t&s() and has not yet written to InsertDone at $e$, if such $tk^{\prime}$ and $ins^{\prime}$ exist. Then $S^{\prime}$ is $S$ followed by $tk_{1},\ldots,tk_{s}$ and $tk^{\prime}$ if it exists, with return value $x_{1},\ldots,x_{s},x^{\prime}$ respectively, in an arbitrary order, with $ins^{\prime}$ placed right before $tk^{\prime}$ if they exist, followed by $tk$ with return value EMPTY if $e$ is a write on LABEL:b-bounded:take:failing_writes_TakeDone (otherwise, $tk$ is one of the $s$ successful Take operations). To show that $S^{\prime}$ is a linearization, we must show that (i) $\{x_{1},\ldots,x_{s}\}\subseteq B$, (ii) $B=\{x_{1},\ldots,x_{s}\}$ if $e$ is a write on LABEL:b-bounded:take:failing_writes_TakeDone, and (iii) $|B|<b$ if $tk^{\prime}$ exists. Due to the no repetitions assumption, (ii) would imply $|B|=s$.
For each $1\leq i\leq s$, let $ins_{i}$ be the Insert operation that performed the last write to Items[$a_{i}$] before $tk_{i}$ read it for the last time (which it did before executing its successful t&s()). Hence, $ins_{i}$ in an Insert($x_{i}$) operation. We begin with proving (i). Let $i$ be an index such that $1\leq i\leq s$. $ins_{i}$ wrote to InsertDone on LABEL:b-bounded:insert:write_InsertDone in $\sigma$ by definition of $tk_{i}$. Thus, $ins_{i}$ is linearized in $\sigma$, namely, it appears in $S$. Since $S$ is a linearization, in the prefix of $S$ that ends with $ins_{i}$, the bag contains $x_{i}$. By Lemma 21, there is no Take operation after $ins_{i}$ in $S$ that takes $x_{i}$. This implies $x_{i}\in B$.
We proceed to prove (ii). We assume $e$ is a write on LABEL:b-bounded:take:failing_writes_TakeDone in this case, hence, $tk$ is a failing Take operation. We will prove that no items other than $x_{i}$ for $1\leq i\leq s$ are in $B$.
Let $C$ be the configuration at the end of $\sigma$ (immediately before $e$), let $\widehat{C}$ be the configuration immediately after $tk$ obtained false from InsertDone.dRead() on LABEL:b-bounded:take:reread_InsertDone, let $C^{*}$ be the configuration immediately after the last previous read of InsertDone by $tk$ on either LABEL:b-bounded:take:read_InsertDone or LABEL:b-bounded:take:reread_InsertDone, let $S^{*}$ be the linearization of the prefix of $\sigma$ ending with $C^{*}$, and let $B^{*}$ be the set of items in the bag at the end of $S^{*}$.
Considering $tk$ obtained false from InsertDone.dRead() right before $\widehat{C}$, the producer does not write to InsertDone between $C^{*}$ and $\widehat{C}$. No write to InsertDone is performed by an uncoupled Insert operation between $\widehat{C}$ and $C$, because otherwise $tk$ would have been linearized at that write.
Suppose there was a successful Insert operation that is in $S$, but not in $S^{*}$, which means it is linearized between configurations $C^{*}$ and $C$. Since no uncoupled Insert operation wrote to InsertDone on LABEL:b-bounded:insert:write_InsertDone between $C^{*}$ and $C$, this Insert operation was coupled with a Take operation. Hence it was immediately taken from the bag after it was inserted into the bag. Therefore, any item in $B$ was inserted to the bag in $S^{*}$ and was in $B^{*}$.
So, suppose there is $x^{*}\in B^{*}$ such that $x^{*}\neq x_{i}$ for all $1\leq i\leq s$. Items inserted by coupled Insert operations are immediately taken from the bag by a Take operation linearized at the same step, so they are not in the bag in the end of the linearization of any prefix of the execution. Hence they are not in $B^{*}$. Therefore, the item $x^{*}$ was inserted into the bag by an uncoupled Insert operation, $ins^{*}$. (Each item in the bag may be uniquely associated with the Insert operation that inserted it thanks to our no repetitions assumption.) This operation wrote $x^{*}$ to Items in some location, $a^{*}$, on LABEL:b-bounded:insert:write_item, and then wrote to InsertDone on LABEL:b-bounded:insert:write_InsertDone before $C^{*}$. We will consider each of the following cases separately: (1) $a^{*}=a_{i}$ for some $1\leq i\leq s$, (2) $a^{*}=a^{\prime}$ (relevant only when $ins^{\prime}$ exists), and (3) $a^{*}\neq a_{i}$ for all $1\leq i\leq s$ and $a^{*}\neq a^{\prime}$. For each case, we will show that a Take operation returning $x^{*}$ appears in $S$ after $ins^{*}$.
We start with case (1). $tk_{i}$ read $x_{i}\neq x^{*}$ from Items[$a^{*}$] on LABEL:b-bounded:take:read_item. We showed in the proof of (i) that $ins_{i}$ appears in $S$. If $ins_{i}$ appears in $S$ before $ins^{*}$, then by Lemma 18, a Take operation that performed a successful TS[$a^{*}$].t&s() appears in $S$ after $ins_{i}$. But this is impossible: in the end of the above proof of (i), we showed that no such Take operations appear in $S$ after $ins_{i}$. Therefore, $ins^{*}$ must appear before $ins_{i}$ in $S$. By Lemma 18, a Take operation that takes $x^{*}$ appears in $S$ after $ins^{*}$, and we are done with case (1).
For case (2), $ins^{*}\neq ins^{\prime}$ since the first appears in $S$ and the latter does not. By Lemma 18, a Take operation that takes $x^{*}$ appears in $S$ after $ins^{*}$.
We proceed to case (3). When $ins^{*}$ wrote to Items[$a^{*}$], TS[$a^{*}$] = 0 by Lemma 13. We consider several sub-cases.
(3a) If $a^{*}$ was removed from Allocated on LABEL:b-bounded:insert:add_to_Allocated before $tk$ read Allocated on LABEL:b-bounded:take:read_Allocated after $C^{*}$, then before this read the producer has removed $a^{*}$ from allocalloc on LABEL:b-bounded:insert:remove_from_alloc, which means the producer passed LABEL:b-bounded:insert:check_TS after $ins^{*}$ was completed. (3b) If $tk$ read $a^{*}$ from Allocated on LABEL:b-bounded:take:read_Allocated after $C^{*}$ and then read $\bot$ from Items[$a^{*}$] on LABEL:b-bounded:take:read_item, then the producer wrote $\bot$ to Items[$a^{*}$] on LABEL:b-bounded:insert:clear_item after $ins^{*}$ was completed, and before this write the producer passed LABEL:b-bounded:insert:check_TS. So in both cases—(3a) and (3b), the producer read 1 from TS[$a^{*}$] on LABEL:b-bounded:insert:check_TS after $ins^{*}$ was completed. Since TS[$a^{*}$] = 0 when $ins^{*}$ wrote to Items[$a^{*}$], some Take operation performed a successful t&s() on TS[$a^{*}$] after $ins^{*}$ wrote to Items[$a^{*}$] and before $\widehat{C}$. Let $tk^{*}$ be the first Take to perform such a t&s(). It appears in $S$ because it is linearized at the latest at the read of 1 from TS[$a^{*}$] by the producer.
(3c) Otherwise, after $C^{*}$, $tk$ read $a^{*}$ from Allocated and then read a non-$\bot$ value from Items[$a^{*}$]. In this case, $tk$ must have performed an unsuccessful t&s() of TS[$a^{*}$] on LABEL:b-bounded:take:t&s. As TS[$a^{*}$] = 0 when $ins^{*}$ wrote to Items[$a^{*}$], another Take operation performed a successful TS[$a^{*}$].t&s() after $ins^{*}$ wrote to Items[$a^{*}$] and before the unsuccessful t&s() by $tk$. Let $tk^{*}$ be the first Take to perform such a t&s(). It appears in $S$ because otherwise it would be one of $tk_{i}$ for some $1\leq i\leq s$, but it is not as it operates on location $a^{*}\neq a_{i}$ for all $1\leq i\leq s$.
In all cases—(3a), (3b) and (3c), $tk^{*}$ returns $x^{*}$ Lemma 17. Additionally, $tk^{*}$ appears in $S$ after $ins^{*}$: $tk^{*}$ is linearized after $ins^{*}$ wrote to Items[$a^{*}$] because $tk^{*}$ performs its successful TS[$a^{*}$].t&s() after that write. Since $ins^{*}$ is uncoupled, $tk^{*}$ cannot be linearized before the write to InsertDone by $ins^{*}$, which is the linearization point of $ins^{*}$.
We proceed to prove (iii). Assume $tk^{\prime}$ exists. Let $C$ be the configuration that appears in $\sigma$ right before the addition of $a^{\prime}$ to alloc by $ins^{\prime}$ on LABEL:b-bounded:insert:add_to_alloc. Let $B_{C}$ be the set of items in the bag at the end of the linearization of $\sigma$’s prefix that ends with $C$. By Lemma 23, $|B_{C}|$ is less than or equal to the number of non-$\bot$ items that are in Items at $C$. By Lemma 10, the locations in Items with non-$\bot$ values at $C$ are a subset of the locations in alloc at the same configuration, and there are less than $b$ locations in alloc according to the check on LABEL:b-bounded:insert:alloc<b. Hence, $|B_{C}|<b$. No Insert operations are linearized in $\sigma$ after $C$, since the producer is executing the coupled Insert $ins^{\prime}$ in this interval, which is linearized only at $e$ that comes right after $\sigma$. Therefore, for any prefix of $\sigma$ that extends the prefix ending with $C$, the bag in the end of the linearization of this prefix does not have more items than $B_{C}$, and we are done.
7.2 Lock-Freedom.
When a Take operation finishes an iteration of the repeat loop and is about to begin another iteration, some Insert operation wrote to InsertDone on LABEL:b-bounded:insert:write_InsertDone between the Take’s penultimate read of InsertDone on LABEL:b-bounded:take:read_InsertDone or LABEL:b-bounded:take:reread_InsertDone and its last read of InsertDone on LABEL:b-bounded:take:reread_InsertDone. This write step completes the Insert operation. When an Insert operation finishes an iteration of the repeat loop and is about to begin another iteration, some Take operation wrote to TakeDone on LABEL:b-bounded:take:successful_writes_TakeDone or LABEL:b-bounded:take:failing_writes_TakeDone between the Insert’s penultimate read of TakeDone on Line 33 or LABEL:b-bounded:insert:reread_TakeDone and its last read of TakeDone on LABEL:b-bounded:insert:reread_TakeDone. This write step completes the Take operation.
8 Discussion
This paper explores strongly-linearizable implementations of bags from interfering objects. We presented the first lock-free, strongly-linearizable implementation of a bag from interfering objects, which is, interestingly, also a linearizable implementation of a queue, but not a strongly-linearizable implementation. We also presented several implementations of bounded bags with a single producer from interfering objects: a wait-free, linearizable implementation of a 1-bounded bag, a lock-free, strongly-linearizable implementation of a 1-bounded bag, and a lock-free, strongly-linearizable implementation of a $b$-bounded bag for any value of $b$.
A direct extension of this work is investigating whether it is possible to extend our bounded bag implementations to support two producers or multiple producers but only one or two consumers. Other open questions are whether there are wait-free, strongly-linearizable implementations of ABA-detecting registers and bags from interfering objects. It would also be interesting to construct a lock-free, strongly-linearizable implementation of a bag from interfering objects, such that, at every point in the execution, the number of interfering objects used is bounded above by a function of the number of processes and the number of elements the bag contains. More broadly, exploring strongly-linearizable implementations of objects beyond bags using interfering objects is a compelling direction for future work.
Acknowledgments
This work was supported in part by the Natural Science and Engineering Research Council of Canada (NSERC) grant RGPIN-2020-04178.